### v1

# Optimizing KV Cache Eviction in LLMs: Adaptive Allocation for Enhanced Budget Utilization

Abstract
Large Language Models have excelled in various fields but encounter efficiency limitations due to the extensive KV cache required for long sequences inference. Many efforts try to evict non-critical cache elements during runtime, thereby reducing cache size within a given memory budget while preserving generation quality. Our reexamination of their underlying principles discerns that prevailing strategies essentially aim to minimize an upper bound of eviction loss within a specific budget allocation. However, we observe that the current practice of uniformly allocating budgets across different attention heads during the eviction procedure tends to degrade the quality of generation posten-eviction. In light of these findings, we propose a simple yet effective adaptive allocation algorithm that not only theoretically ensures its loss upper bound does not exceed that of previous uniform allocation methods, but also effectively aligns with the characteristics of the self-attention mechanism, thus practically reducing the upper bound. Further, integrating this algorithm with two of the most advanced methods yields Ada-SnapKV and Ada-Pyramid. Extensive experimental validation across 16 datasets and the Needle-in-a-Haystack test confirm that Ada-SnapKV and Ada-Pyramid achieve further enhancements, establishing new benchmarks in state-of-the-art performance.
[…]
1 Introduction
Autoregressive Large language models (LLMs) have achieved significant success and are widely utilized across diverse natural language processing applications, including dialogue systems (Yi et al. 2024), document summarization (Laban et al. 2023), and code generation (Gu 2023). Their deployment in real-world contexts has propelled the development of their capacity to process extended sequences. For instance, GPT-4 supports sequences up to 129K (Achiam et al. 2023), Claude3 up to 200K (Anthropic 2024), and Gemini-Pro-1.5 (Reid et al. 2024) up to 1M tokens. However, the computational cost and memory requirements for Transformer architecture increase quadratically with sequence length, posing significant challenges.
To mitigate these issues, a traditional optimization technique called caching has been applied to LLMs, referred to as Key-Value(KV) cache. The KV cache pairs store intermediate computation results of previous tokens, effectively reducing the computational burden associated with autoregressive generation. It has become a prevalent approach in enhancing LLM efficiency. However, the expansion of the inference sequence necessitates a corresponding increase in KV cache size, which can easily exceed the total parameter size of the model itself in long-sequence inference  (Sun et al. 2024). This not only leads to substantial memory overhead but also results in massive input/output (I/O) time during autoregressive generation (Sun et al. 2024).
In response, KV cache eviction methods (Zhang et al. 2024b; Xiao et al. 2023; Ge et al. 2023; Yang et al. 2024; Zhang et al. 2024a; Li et al. 2024) have been proposed to compress the cache size within specified budgets by discarding non-critical cache pairs while striving to minimize the loss of generation quality. These methods are valued for their plug-and-play capabilities, allowing for seamless integration into any LLM without fine-tuning. Typically, these algorithms employ various strategies to select and evict the majority of KV cache pairs, thus reducing memory demands. The latest leading algorithms use the Top-K base selection scheme, which effectively distinguishes between critical and non-critical cache pairs, retaining the former and evicting the latter. The recent leading algorithms have established the Top-K base selection scheme, which can effectively to distinguish the critical and non-critical cache pairs thus perserving and evicting. Despite these advancements, the challenge of minimizing quality loss in existing eviction schemes remains unresolved.
Our study begins by reexamining the underlying principles of eviction strategies from a theoretical perspective. We establish that Top-K based strategies aim to minimize an upper bound of eviction loss, quantified by the L1-distance between the outputs of the self-attention mechanism pre- and post-eviction under given budget allocation results. We also found the common practice of distributing budget uniformly across different attention heads leads to a misallocation of total budgets, thus diminishing the generation quality, as illustrated in Figure 1(a). Base on these insights, we propose a simple yet effective adaptive allocation algorithm to optimize budget distribution to improve the generation quality, which utilizes the pronounced variations in attention concentration among the heads within the self-attention mechanism as illustrated in Figure 1(b). In addition, we also theoretically prove that the upper bound of the loss under Top-K eviction with adaptive allocation consistently remains at or lower than that with uniform allocation, ensuring the robustness of algorithm performance in practice.
By integrating the adaptive budget allocation algorithm into two prominent methods (Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a), we develop two adaptive eviction methods: Ada-SnapKV and Ada-Pyramid, respectively. Extensive evaluations across 16 datasets spanning various tasks in LongBench demonstrate that both Ada-SnapKV and Ada-Pyramid further improve the generation quality. Another widely used test, named ”Needle-in-a-Haystack”, also shows that adaptive allocation further enhances long-context retrieval capabilities. The main contributions are summarized as follow:
  • Reexamination of Cache Eviction Principle: We formalize the principle of cache eviction as the minimization of the upper bound of eviction loss, quantified as changes in the L1 distance of attention outputs post-eviction.
  • Identifying and Addressing Limitations in Budget Allocations: We pinpoint the inefficiency of existing uniform budget allocation across heads. To address this, we introduce an adaptive allocation algorithm that both theoretically and empirically reduces the upper bound of eviction loss.
  • Developments of Adaptive Eviction Methods: Base on the proposed adaptive allocation algorithm, we develop two novel approaches, Ada-SnapKV and Ada-Pyramid. These methods demonstrably mitigate quality loss during compression in the extensive evaluations.
  • Innovative Direction to Enhance Cache Eviction: Our research presents a method that adaptively allocates budgets based on unique characteristics among heads to optimize cache eviction, guided by our theoretical analysis of eviction loss upper bounds. We hope this direction achieves significant further development.
(a) Uniform Allocation
(b) Adaptive Allocation
Figure 1: An Illustrative Example.This example features a five KV cache pairs with associated attention weights across three heads. Adaptive allocation, through reallocating budgets from the concentrated Head2 to the dispersed Head1, increases the aggregated weights of retained KV cache pairs from 2.33 to 2.46, which correlates to a reduced upper bound of eviction loss in Section 3.5.
2 Related Works
In the context of long sequences, the vast scale of the KV caches leads to a memory-bound situation, causing significant memory and I/O latency costs  (Wang and Chen 2023). Some studies, which are orthogonal to our work, employ efficient memory management strategies that reduce I/O time without altering the size of the KV Cache, such as Page Attention (Kwon et al. 2023) and Flash Attention (Dao et al. 2022). And in our experiments, we have incorporated the Flash Attention technique to achieve efficient computation. Other approaches (Tang et al. 2024) attempt to reduce I/O overhead by only recalling KV cache entries relevant to the current query for computation, while others remain stored for subsequent queries. However, these methods are constrained by substantial memory overheads, making it difficult to deploy on GPUs with lower storage capacities.
Recently, KV cache eviction methods have gained attention and rapid development due to their flexible compression ratios given any storage budgets and the advantage of being plug-and-play without the need for fine-tuning. Earlier StreamingLLM (Xiao et al. 2023) simply maintains the cache of 4 initial and the recent tokes, discarding all others to adapt to long sequence inference. FastGen (Ge et al. 2023) searches and combines multiple strategies, like maintaining the cache of special tokens, punctuation tokens, and recent tokens, based on the characteristics of attention heads. H2O (2024b) has developed an eviction algorithm that utilizes query states of all tokens to identify important KV pairs based on the Heavy Hitter method. The most recent SnapKV (Li et al. 2024) identifies important KV cache pairs using query states of several tokens within a recent window, evicting the less important ones. This method effectively mitigates the quality degradation in cache eviction. The Pyramid (Yang et al. 2024; Zhang et al. 2024a) further adjusts the budget allocation across different layers in SnapKV, improving the generation quality in small-budget scenarios. However, to our best knowledge, current eviction methods have never tried to adaptively distribute the total budget across different heads. Building on theoretical analysis and observations of inherent attention patterns in LLM heads, we have identify and demonstrate the necessity for adaptive budget allocation in cache eviction. Based on these insights, we propose a simple yet effective budget allocation mechanism and integrate it into the two leading strategies, SnapKV and Pyramid, further reducing the accuracy drop associated with cache eviction.
3 Framework
3.1 Overview
In this section, we reexamine how the existing cache eviction strategies to retain essential information in the past KV cache from a theoretical perspective. Inspired by theoretical findings , we propose a simple yet effective algorithm for adaptive budget allocation, which is proved to better than the previous uniform allocation in cache eviction procedure both in theory and practice. Further integrating it into two current leading methods, we develop two adaptive cache eviction methods: Ada-SnapKV and Ada-Pyramid. The key findings and insights are as below:
  • How the eviction strategies maintain generation quality despite discarding substantial KV cache? Theorem 2 formalizes an upper bound of the eviction loss by using the L1-distance between the before and post-eviction outputs and Theorem 3 demonstrates that Top-K based eviction strategies actually minimize this upper bound.
  • Why is adaptive budget allocation better from the theoretical perspective? Theorem 4 demonstrates that the upper bound of eviction loss under adaptive allocation maintains at or below that under previous uniform allocation in the cache eviction procedure. Thus it implies that the adaptive allocation is better in theory.
  • Why is adaptive budget allocation better from the empirical perspective? As shown in Figure 2 and 3, we identify that the inherent disparities in attention concentration across different heads reliably translate theoretical advantages into practical results. This leads to consistent reductions of eviction loss upper bound across each layer in eviction procedure.
3.2 Preliminary
We begin by providing a formal description of the computational processes involving KV cache and Multi-head Attention in a single-layer of LLMs to alleviate the burden of notation. LLMs are characterized by their autoregressive generation mode, where each step involves using the current final token to predict the next token. Define $X\in\mathbb{R}^{n\times d}$ as the embedding matrix of all tokens in sequence, and $x\in\mathbb{R}^{1\times d}$ as the last token used as input at the current timestep. To clarify the subsequent theoretical exposition, we adopt the notation system from  (Liu et al. 2023) under the assumption of $h$ attention heads in one layer. The transformation matrices for each head $i\in[h]$ map token embeddings to their respective Query, Key, and Value are denoted as $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\,\in\mathbb{R}^{d\times d_{h}}$ , and the final output matrix $W_{i}^{O}\,\in\mathbb{R}^{d_{h}\times d}$ transform the intermediate result to the output hidden states. At each timestep, the states of the stored KV cache for head $i$ has been initialized as:
$K_{i}=XW_{i}^{K},V_{i}=XW_{i}^{V}$ (1)
Then, the input token $x$ is mapped to its corresponding query, key,value for each head, and the previous KV cache is updated accordingly:
$q_{i}=xW_{i}^{Q},k_{i}=xW_{i}^{K},v_{i}=xW_{i}^{V}$ (2)
$K_{i}=[K_{i}:k_{i}],V_{i}=[V_{i}:v_{i}]$ (3)
Finally the final output $o\in\mathcal{R}^{1\times d}$ is computed as follow:
$o=\sum^{i\in[h]}A_{i}V_{i}W_{i}^{O}$ (4)
where $A_{i}\in\mathbb{R}^{1\times n}$ is the attention weight calculated by:
$A_{i}=\text{softmax}(q_{i}K_{i}^{T})$ (5)
3.3 Reexamining the Principle of Cache Eviction
Cache eviction is dedicated to reducing the size of KV cache to fit within a constrained budget by evicting certain cache pairs strategically. Eviction masks $\left\{\mathcal{M}_{i}\in\mathbb{R}^{1\times n}\right\}$ can be employed to simulate the post-eviction output $o^{\prime}$ in self-attention mechanism:
$o^{\prime}=\sum^{i\in[h]}A_{i}^{\prime}V_{i}W_{i}^{O}\>\text{where}\>A_{i}^{\prime}=\text{softmax}(q_{i}K_{i}^{T}+\mathcal{M}_{i})$ (6)
where each element $\mathcal{M}_{i}^{j}$  [^1] Given the first dimension of $\mathcal{M}_{i}$ is 1, $M_{i}^{j}$ is used to simplify the notation for $\mathcal{M}_{i}(1,j)$. Similarly, $A_{i}^{j}$ is in the same manner. in mask $\mathcal{M}_{i}$ indicates that whether evicting the $j\text{th}\in[n]$ KV cache pair in $K_{i},V_{i}\in\mathbb{R}^{n\times d_{h}}$ on each head $i$:
$\mathcal{M}_{i}^{j}$ =$=\begin{cases}0&\text{if the $j$th cache pair on head $i$ is retained }\\ -\infty&\text{otherwise the $j$th cache pair on head $i$ is evicted}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
Thus, the budget $B_{i}$ for head $i$ corresponds to the number of zero elements in $\mathcal{M}_{i}$. Theorem 1 further simplify the output $o^{\prime}$ by eliminating the softmax function. A detailed proof is provided in Appendix A.1.
Theorem 1.
The post-eviction output $o^{\prime}$ can rewrite as:
$o^{\prime}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (7)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
The reduction in generation quality due to cache eviction stems from alterations of the attention output. Thus, we model the eviction loss using the L1-distance between the post-eviction and original outputs of self-attention mechanism:
Eviction $\text{Eviction Loss}=||o^{\prime}-o||_{1}$ (8)
Utilizing the L1-norm of matrix, which can understand how the matrix transforms vectors, we derive an upper bound $D$ for the Eviction Loss in Theorem 2. For a detailed proof, refer to Appendix A.2.
Theorem 2.
The eviction loss caused by cache eviction can be bounded by $D$ as follows:
Eviction $\text{Eviction Loss}\leq D=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (9)
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
where $C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{\infty}\right\}$ is the max value in the row norms of Matrices $\left\{V_{i}W_{i}^{O}\right\}$ among all heads.
As shown in Algorithm 1, the core idea of current leading eviction strategies based on Top-K selections is to retain part of cache pairs corresponding to $B_{i}$ highest weights $A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})$, while evicting those considered non-essential. Obviously, given the budget allocation $\left\{B_{i}\right\}$ the Top-K based eviction strategies maximize the aggregated weights as follow:
Top-K Eviction $=\operatorname*{arg\,max}_{strategy}\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}.$ (10)
[…]
Therefore, we have established the following Theorem 3.
Algorithm 1 Eviction Based on Top-K Selection
Input: Allocation $\left\{B_{i}\right\}$,KV Cache $\left\{K_{i},V_{i}\right\}$, Weights $\left\{A_{i}\right\}$
Output: Compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
1:  for $i\leftarrow 1$ to $h$ do
2:     initialize empty cache $\hat{K}_{i},\hat{V}_{i}$ for head $i$
3:     for $j\leftarrow 1$ to $n$ do
4:        if $A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})$ then
5:           retain and append cache pair $K_{i}^{j},V_{i}^{j}$ to $\hat{K}_{i},\hat{V}_{i}$
6:        else
7:           evict cache pair $K_{i}^{j},V_{i}^{j}$
8:        end if
9:     end for
10:  end for
11:  return  compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
Theorem 3.
Given a budget allocation $\left\{B_{i}\right\}$, the principle of eviction strategies based on Top-K selections is to minimizes the upper bound $D$ of the eviction loss.
[…]
$\>\text{given}\>\left\{B_{i}\right\}\>\text{s.t.}\>\sum^{i\in[h]}B_{i}=B$
3.4 Adaptive vs. Uniform: Theoretical Insights.
In existing studies on cache eviction, the budget $B$ is uniformly distributed across each head within a layer: specifically, ${B_{i}=B/h}$. Consequently, the upper bound of eviction loss $D$ under any given allocation, is modified to $D^{\prime}$ under uniform allocation:
$D^{\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}$)}}A_{i}^{j}$ (12)
given uniform $\text{given uniform allocation}\left\{B_{i}=B/h\right\}$
In contrast, we suggest adaptive distribution of the total budget among different heads and introduce a simple yet effective budget allocation algorithm that dynamically distributes the total budget $B$ based on the attention weights $A_{i}$ across $h$ heads, with the allocated results defined as $\left\{B^{*}_{1},B^{*}_{2},...,B^{*}_{h}\right\}$ subject to $\sum^{i\in[h]}B^{*}_{i}=B$. As shown in Algorithm 2, it firstly selects the $B$ largest attention weights from all heads within one layer. Based on the times of each head is selected in the above procedure, different budgets ${B^{*}_{i}}$ are allocated to each head. Under this adaptive allocation, the upper bound of eviction loss is denoted as $D^{\prime\prime}$:
$D^{\prime\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}^{*}$)}}A_{i}^{j}$ (13)
given adaptive $\text{given adaptive allocation}\left\{B_{i}^{*}\right\}$
Theorem 4.
The upper bound $D^{\prime\prime}$ of eviction loss with adaptive budget allocation consistently remains at or below the upper bound $D^{\prime}$ associated with uniform allocation.
$D^{\prime\prime}\leq D^{\prime}$ (14)
Algorithm 2 Adaptive Budget Allocation
Input: Total Budget $B$, Attention Weights in $h$ heads $\left\{A_{i}\in\mathbb{R}^{1\times n}\right\}$;
Output: Allocated Budgets of $h$ heads $\left\{B^{*}_{i}\right\}$;
1:  Concatenate across heads $A$ = Cat($\left\{A_{i}\right\}$,dim=1)
[…]
3:  Identify top indices $T$ = Top-K($A,B$).indices
4:  Select the corresponding head indicator $I^{*}=I[T]$
5:  Count frequencies of each $i$ in $I^{*}$ to determine $\left\{B^{*}_{i}\right\}$
6:  Return Allocated Results of $h$ heads $\left\{B^{*}_{i}\right\}$
According to Theorem 4, our adaptive allocation algorithm achieves equal or smaller eviction loss than the previous uniform allocation approach, thereby enhancing the post-eviction generation quality. The detailed proof can be found in Appendix A.3.
Figure 2: Disparities in Attention Concentration Across Heads. (Mistral-instruct-v0.2 on the first sample of Qasper, a single-doc QA dataset in LongBench) We aggregate different proportions of the top attention weights, $\sum^{\text{retained}j}A_{i}^{j}$, to analyze attention concentration in different head $i$. Most concentrated heads require a small cache proportion, e.g., $5\%$, to aggregate weights close to 1, whereas other dispersed heads need significantly larger proportions, such as $20\%$.
3.5 Adaptive vs. Uniform: Empirical Insights
According to Theorem 4, the upper bound of eviction loss of adaptive allocation is equal or less than to that of the current uniform allocation. We further demonstrate different attention heads within each layers of LLMs exhibit significant disparities in attention concentration, resulting in the necessity of adaptive budget allocation in practice. For visualization in Figure 2 (a), most concentrated heads in Layer 8 like head 1 require only 1% of the original cache budget to effectively retain the aggregated weights $\sum^{\text{retained}j}A_{i}^{j}$ of 0.95. Conversely, other heads like heads 18 requires nearly 50% proportion to near 0.95. This characteristic is closely related to the upper bound $D$ of eviction loss, as detailed in Theorem 3. Under such circumstances, the previous uniform budget allocation faces a dilemma, as an illustrative example in Figure 1: either neglect the loss in dispersed heads, or allocate excessive and unnecessary budgets to heads with concentrated attention. This significantly undermines the trade-off performance between the total budget and generation quality in previous cache eviction. In contrast, the adaptive allocation algorithm assigns large budgets to dispersed heads, while controlling the budget sizes for other heads, effectively maintaining the overall budget size and mitigating the decline in generation quality. In Figure 3, we also visualize the gap between the upper bounds of eviction loss, $D^{\prime}$ and $D^{\prime\prime}$, under previous uniform allocation and proposed adaptive allocation of an average budget size of 128 for each head. Results show that adaptive allocation consistently reduces eviction loss upper bound, especially in layers 8 to 30.
Figure 3: Comparison of Upper bounds. (In the instance of Mistral-instruct-v0.2 on the first sample of Qasper, a single-document QA dataset in LongBench, adaptive allocation consistently reduced eviction loss upper bound.)
3.6 Implementation
The current two leading eviction strategies, SnapKV and Pyramid, both utilize the several tokens $X^{rec}\in\mathbb{R}^{win*d}$ from a recent window (typically window size is 32) to identify and evict the less important cache pairs. SnapKV excels at managing evictions under scenarios with large budgets, while Pyramid is more effective in environments with smaller budgets. The key distinction lies in how Pyramid and SnapKV allocate budgets within different layers in LLMs. Pyramid suggests that information aggregation among layers takes a pyramidal form, thereby allocating a larger budget to shallower layers and progressively reducing it in deeper layers through pre-set hyper-parameters. In contrast, SnapKV distributes the budget uniformly across all layers. Nonetheless, same as other eviction methods, they both allocate the budget uniformly among all heads within a single layer. We incorporate the adaptive allocation algorithm into both of them, resulting in the creation of two novel adaptive strategies Ada-SnapKV and Ada-Pyramid, respectively.
Take the Ada-SnapKV in algorithm 3 as an example [^2] Algorithm 3 describes the process serially for simplicity, but eviction operations can be parallelized easily in practice., the adaptive budget allocation can be seamlessly integrated into any eviction strategy by invoking Algorithm 2 before eviction process in each layer to distribute adaptively the total budget among all heads. In line 6, a hyper-parameter $\alpha$, set by default to 0.5, prevent from assigning tiny budgets to highly concentrated heads, enhances fault tolerance in the post-eviction generation.
Algorithm 3 Ada-SnapKV in One Layer
Input: Total budget $B$, Past cache $\left\{K_{i},V_{i}\right\}$, Tokens in the recent window $X^{rec}\in\mathbb{R}^{win*d}$
Output: Compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
1:  for $i\leftarrow 1$ to $h$ do
2:     $Q_{i}^{rec}=X^{rec}W_{i}^{Q}$
3:     $\bar{A_{i}}=softmax(Q_{i}^{rec}K_{i}^{T})$
4:     $\bar{A_{i}}=\bar{A_{i}}.maxpooling(dim=1).mean(dim=0)$
5:  end for
6:  get $\left\{B_{i}^{*}\right\}$ by invoking Algorithm 2($B,\left\{\bar{A}_{i}\right\}$)
7:  $\left\{B_{i}^{*}\right\}=\alpha\times\left\{B_{i}^{*}\right\}+(1-\alpha)\times\frac{B}{h}$
8:  $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$ = Algorithm 1($\left\{B_{i}^{*}\right\},\left\{K_{i},V_{i}\right\},\left\{\bar{A}_{i}\right\}$)
9:  return  compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
Single-Doc. QA Multi-Doc. QA Summarization Few-shotLearning Synthetic Code
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Ave. Score
Full Cache 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 55.33 52.87   42.51
B=128h
H2O 21.19 21.66 38.60 30.63 20.65 12.19 20.65 22.42 21.81 39.00 82.52 40.68 2.98 79.56 49.13 46.76  34.40
StreamingLLM 16.61 14.74 31.40 28.05 21.36 12.08 18.44 18.91 19.26 43.50 74.22 29.00 2.75 31.65 41.27 38.84  27.63
SnapKV 19.17 21.40 42.93 36.76 22.44 15.86 19.16 21.84 21.55 47.50 84.15 40.24 2.30 68.26 50.69 47.13  35.09
Pyramid 20.16 21.77 43.55 36.78 23.12 14.39 19.53 22.03 21.47 51.00 84.62 40.24 2.79 70.77 50.57 46.53  35.58
Ada-SnapKV 20.63 22.58 45.68 37.90 23.49 16.55 19.99 22.28 21.55 59.50 85.00 40.62 3.09 69.36 50.98 48.17  36.71
Ada-Pyramid 20.50 21.71 45.61 36.81 23.57 15.84 19.75 22.13 22.00 60.50 84.04 40.51 3.21 73.60 51.24 48.02 36.81
B=256h
H2O 21.54 22.92 42.56 31.07 22.53 13.76 22.52 22.40 23.09 40.50 84.20 40.77 3.41 86.10 50.98 48.17  36.03
StreamingLLM 17.93 16.01 33.36 30.71 21.30 10.08 20.66 19.47 22.89 53.50 73.59 29.22 3.00 27.77 42.30 39.87  28.85
SnapKV 22.37 23.74 48.13 38.56 22.43 15.66 21.91 23.13 23.15 61.50 85.45 41.42 3.09 84.54 53.22 50.24  38.66
Pyramid 20.09 24.00 47.33 38.24 22.48 16.02 21.40 22.45 22.63 63.00 84.93 40.98 3.40 82.48 52.78 49.36  38.22
Ada-SnapKV 22.55 25.78 48.33 40.30 24.24 16.64 21.63 23.03 23.19 67.00 85.78 41.53 3.47 87.07 53.86 51.13 39.72
Ada-Pyramid 22.64 24.64 47.40 40.25 23.62 16.83 21.82 23.34 22.70 66.50 84.99 41.34 2.78 86.90 53.17 49.52  39.28
B=512h
H2O 21.72 26.03 44.81 32.33 23.16 14.86 23.65 22.84 24.70 42.00 85.22 41.57 3.40 86.45 53.04 49.68  37.22
StreamingLLM 18.76 17.17 37.09 30.21 21.64 9.93 24.44 20.00 25.57 62.00 72.36 29.95 2.48 18.17 43.70 40.13  29.60
SnapKV 24.60 27.81 48.98 39.46 25.25 16.98 23.70 22.96 24.37 67.00 85.88 41.26 2.78 86.56 54.81 51.71  40.26
Pyramid 23.23 27.94 48.87 40.50 24.36 16.74 23.22 23.16 24.37 67.00 85.73 41.74 3.16 85.67 54.16 50.34  40.01
Ada-SnapKV 23.39 28.72 48.96 40.60 25.20 17.25 23.15 23.48 24.41 68.00 86.39 41.69 2.73 88.92 54.69 51.51 40.57
Ada-Pyramid 24.03 28.98 48.39 39.25 24.50 18.38 23.13 23.90 24.30 68.00 85.89 41.89 2.98 87.71 54.46 51.39  40.45
B=1024h
H2O 23.90 28.62 46.46 37.03 24.74 15.04 25.30 23.11 25.92 46.00 85.93 41.80 3.24 86.57 54.46 51.01  38.70
StreamingLLM 19.42 21.69 41.75 32.40 22.18 11.18 27.13 21.09 26.59 67.00 71.79 30.11 2.88 16.57 44.82 39.76  31.02
SnapKV 25.47 29.57 49.33 40.90 25.53 19.01 25.94 23.89 26.21 69.50 86.48 42.10 2.98 88.56 55.57 51.92  41.44
Pyramid 24.21 29.86 48.93 40.75 25.05 18.77 25.73 24.06 25.65 68.50 86.31 42.25 2.97 87.17 54.75 52.10  41.07
Ada-SnapKV 24.79 31.94 48.45 40.73 26.22 19.11 25.61 23.92 26.03 70.00 86.32 42.35 2.91 88.31 55.44 52.55 41.54
Ada-Pyramid 25.09 30.94 48.18 40.00 26.52 19.10 24.93 23.71 25.86 70.00 86.34 42.64 2.56 86.92 54.93 51.90  41.23
Table 1: Comparison Based on Mistral-7B-Instruct-v0.2 Among 16 Datasets
4 Experiments
4.1 Settings
Datasets
Firstly, we carry out an comprehensive evaluation using 16 datasets, covering domains of single-document QA (Kočiskỳ et al. 2018; Dasigi et al. 2021), multi-document QA (Yang et al. 2018; Ho et al. 2020; Trivedi et al. 2022), summarization (Huang et al. 2021; Zhong et al. 2021; Fabbri et al. 2019), few-shot learning (Joshi et al. 2017; Gliwa et al. 2019; Li and Roth 2002), Synthetic (Bai et al. 2023), and code generation (Guo et al. 2023; Liu, Xu, and McAuley 2023) , within LongBench (Bai et al. 2023), a benchmark for evaluating multi-task performance with long-sequence inputs. These datasets feature varying average input lengths from 1,235 to 18,409 tokens, necessitating substantial KV cache size during generation, thereby rendering them suitable for evaluating KV cache eviction strategies. Each dataset is assessed using LongBench-recommended metrics, with the quality scores up to 100. Detailed dataset information is provided in Appendix A.5. We also utilize the widely-used ’Needle-in-a-Haystack’ test, where key information is randomly inserted into long texts to create prompts. This test evaluates whether LLMs can extract this key information from extensive texts, specifically examining the impact of proposed adaptive allocation on the models’ fundamental long context retrieval abilities.
(a) SnapKV
(b) Pyramid
(c) Ada-SnapKV
(d) Ada-Pyramid
Figure 4: Needle-in-a-Haystack Test. This test inserts a critical sentence (the ’needle’) within a document (the ’haystack’) with extensive context, then evaluates a model’s ability to retrieve the needle from the document. The x-axis indicates the context length of the document, and the y-axis shows the insertion depth of the needle. The Average Score is determined by averaging the aggregated scores at various context lengths. Higher scores indicate an improved capacity of the model for contextual retrieval.
Baselines
We conduct a comparative analysis between our newly proposed methods, Ada-SnapKV and Ada-Pyramid, and the previously established SnapKV and Pyramid, acknowledged as two leading approaches under varying scenarios. Additionally, we assesse these methods against other earlier eviction strategies, including StreamingLLM and H2O. StreamingLLM is designed to handle long-sequence inputs by retaining caches of several initial tokens and others in a recent window. And H2O compresses cache size by employing attention weight-based detection of Heavy Hitters based on all query and key states. In all experiments, the hyper-parameter $\alpha$ in Algorithm 2 is set to 0.5. Both Ada-SnapKV and Ada-Pyramid, as well as SnapKV and Pyramid, utilize the same configuration settings as described in (Li et al. 2024), ensuring comparability with a recent window size of 32 and a maximum pooling kernel size of 7. Parameters for StreamingLLM and H2O conform to the default settings reported in the literatures (Zhang et al. 2024b; Xiao et al. 2023).
Base Models
In the experiments, we employ two open-source base models: Mistral-7B-instruct-v0.2 (Jiang et al. 2023) and LWM-Text-Chat-1M (Liu et al. 2024). The Mistral 7B model features a context length of 32K and has been adopted as the primary model in related studies (Li et al. 2024; Zhang et al. 2024a) due to its moderate parameter size and remarkable capability for long text tasks. Meanwhile, LWM stands as the state of the art with its 1M context length, facilitating performance evaluations for Needle-in-a-Haystack test under extreme context lengths.
(a) Mistral-7B-Instruct-v0.2
(b) LWM-Text-Chat-1M
Figure 5: Average Score Among 16 Datasets
4.2 Evaluations Among 16 Datasets
We assess all eviction strategies using cache budget $B\in\left\{128\times h,256\times h,512\times h,1024\times h\right\}$ for each layer. Detailed results for each dataset on the Mistral model are provided in Table 1, while other results for the LWM model are placed in Appendix A.4 due to space constraints. To demonstrate the efficacy of adaptive allocation, we take a budget $B=128h$ as an example presented in Table 1. After integrating the adaptive allocation algorithm, Ada-SnapKV enhances the quality scores in 15 out of 16 datasets compared to the original SnapKV, increasing the average score from 35.09 to 36.71. Similarly, Ada-Pyramid surpasses the original Pyramid in 14 of 16 datasets, boosting the average score from 35.58 to 36.81.
Figure 5 summarizes performance based Mistral and LWM across 16 datasets. Overall, SnapKV and Pyramid, as the current two leading methods, exhibit closely matched performance, surpassing previous approaches such as H2O and StreamingLLM. Furthermore, our Ada-SnapKV and Ada-Pyramid strategies consistently improve the generated quality under varying budgets, especially in small budgets. The two adaptive eviction strategies with alternatingly leading and surpassing previous versions to become the new state-of-the-art methods. This consistent improvement validates the necessity and effectiveness of adaptive budget allocation, as demonstrated in both theoretical derivations and empirical findings.
4.3 Evaluations on Needle-in-a-Haystack Test
As shown in Figure 4, we employ a Needle-in-a-Haystack test to demonstrate how adaptive budget allocation can enhance long-context retrieval capabilities. All configurations maintains a recent window size of 32 and a pooling kernel size of 7 which consistent with former experiments, where the maximum inference length is limited to 33K in the full cache case on A100-80G. With a cache budget of $B=128h$, all four strategies—Ada-SnapKV, Ada-Pyramid, SnapKV, and Pyramid—successfully extend inference length up to 217K. Notably, Ada-SnapKV and Ada-Pyramid both effectively improve long-text retrieval capabilities. In particular, Ada-SnapKV and Ada-Pyramid achieve near-lossless retrieval within the original 33K length, a feat not replicated by the standard SnapKV and Pyramid. In terms of average score, Ada-SnapKV improves from 95.40 to 96.56, while Ada-Pyramid increases from 96.62 to 97.02.
5 Conclusion
In this study, we reexamine prevailing cache eviction strategies employed in LLMs, discerning their goal to minimize an upper bound of eviction loss. This loss is quantified as the L1 distance between outputs before and after eviction. By introducing an adaptive budget allocation among various attention heads, we theoretically reduce the upper bound compared to previous practices. Our empirical findings suggest that this adaptive approach significantly benefits from the varied concentration levels inherent among multiple heads within the self-attention mechanism. We develop two novel adaptive eviction methods, Ada-SnapKV and Ada-Pyramid, which incorporate this adaptive allocation into two advanced existing strategies. Comprehensive evaluation confirms that both methods yield substantial improvements, especially in small budgets. A comprehensive evaluation verifies that both methods significantly enhance performance, particularly in scenarios with small budgets. Furthermore, this research emphasizes the significant potential for advancing cache eviction strategies through our theoretical framework and adaptive budgeting, which are specifically designed to exploit the unique characteristics of different attention heads in LLMs.
Appendix A Appendix
A.1 Proof of Theorem 1
Theorem.
The post-eviction output $o^{\prime}$ can rewrite as:
$o^{\prime}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (15)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
Proof.
Consider the softmax function as:
$softmax(x)=\frac{exp(x^{j})}{\sum^{j}exp(x^{j})}$ (16)
Thus, the attention weight after eviction procedure is:
$A^{\prime}_{i}=softmax(s_{i}+\mathcal{M}_{i})\>where\>s_{i}=q_{i}K_{i}^{T}$ (17)
$A^{\prime}_{i}$ $=\frac{exp(s_{i}^{j}+\mathcal{M}_{i}^{j})}{\sum^{j}exp(s_{i}^{j}+\mathcal{M}_{i}^{j})}=\frac{exp(s^{j}_{i})\odot N_{i}^{j}}{\sum^{j}exp(s^{j}_{i})\odot\mathcal{N}_{i}^{j}}$ (18)
$=\frac{exp(s^{j}_{i})\odot N_{i}^{j}}{\sum^{j}exp(s^{j}_{i})}\frac{\sum^{j}exp(s^{j}_{i})}{\sum^{j}exp(s^{j}_{i})\odot\mathcal{N}_{i}^{j}}$ (19)
$=\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}$ (20)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
Thus:
$o^{\prime}=\sum^{i\in[h]}A_{i}^{\prime}V_{i}W_{i}^{O}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (22)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
A.2 Proof of Theorem 2
Theorem.
The eviction loss caused by cache eviction can be bounded by $D$ as follows:
Eviction $\text{Eviction Loss}\leq D=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (23)
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
where $C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{1}\right\}$ is the max value in the L1-norm of Matrices $\left\{V_{i}W_{i}^{O}\right\}$ among all head.
Proof.
By calculating the L1 distance between their outputs, we can obtain
$||o^{\prime}-o||_{1}=||\sum^{i\in[h]}(\textbf{1}-\frac{\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}V_{i}W_{i}^{O}||_{1}$ (24)
$\leq\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}V_{i}W_{i}^{O}||_{1}$ (25)
$\leq\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{\lVert A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}\rVert_{1}\>\lVert V_{i}W_{i}^{O}\rVert_{\infty}$ (26)
$\leq C\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{\lVert A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}\rVert_{1}$ (27)
$where\>C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{\infty}\right\}$
By expanding $A_{i}$, we can further simplify the expression.
$||o^{\prime}-o||_{1}\leq C\sum^{i\in[h]}\sum^{\text{retained}j}(\frac{A_{i}^{j}}{\sum^{\text{retained}j}A_{i}^{j}}-A_{i}^{j})+\sum^{\text{evicted}j}A_{i}^{j}$ (28)
$=C\sum^{i\in[h]}\sum^{\text{retained}j}(\frac{A_{i}^{j}}{\sum^{\text{retained}j}A_{i}^{j}})-\sum^{\text{retained}j}A_{i}^{j}+\sum^{\text{evicted}j}A_{i}^{j}$ (29)
$=C\sum^{i\in[h]}(2-2\sum^{\text{retained}j}A_{i}^{j})$ (30)
$=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (31)
A.3 Proof of Theorem 4
Theorem.
The upper bound $D^{\prime\prime}$ of eviction loss with adaptive budget allocation consistently remains at or below the upper bound $D^{\prime}$ associated with uniform allocation.
$D^{\prime\prime}\leq D^{\prime}$ (32)
Proof.
$D^{\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}$)}}A_{i}^{j}$ (33)
given uniform $\text{given uniform allocation}\left\{B_{i}=B/h\right\}$
$D^{\prime\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}^{*}$)}}A_{i}^{j}$ (34)
given adaptive $\text{given adaptive allocation}\left\{B_{i}^{*}\right\}$
Based on the operations of concatenation and Top-K in the first and second lines of Algorithm 2, we can flatten the second $\sum$ in $D^{\prime\prime}$ as follow:
$D^{\prime\prime}=2hC-2C\sum^{i\in[h],j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A,B)}A_{i}^{j}$ (35)
$\text{where}\>B=\sum^{i}B_{i}^{*}=\sum^{i}B_{i}$
Considering that $B=\sum{B_{i}}=\sum{B_{i}^{*}}$ , it is evident that $\sum^{i\in[h],j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A,B)}A_{i}^{j}\geq\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})}A_{i}^{j}$. This is because under the premise of identical total budget, the global Top-K sum is greater than or equal to the sum of local Top-K sums of each head. Thus:
$D^{\prime\prime}\leq D^{\prime}$ (36)
Figure 6: Visualization of Heads’ Concentrations
Single-Doc. QA Multi-Doc. QA Summarization Few-shotLearning Synthetic Code
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Ave. Score
Full Cache 18.00 25.80 43.10 23.40 16.70 9.70 27.20 25.00 24.70 70.50 61.60 39.60 3.00 6.50 42.20 41.60  29.91
B=128h
H2O 17.90 17.73 36.10 21.52 17.51 9.26 16.13 22.99 19.64 43.50 60.64 36.36 3.00 5.50 34.93 36.74  24.97
StreamingLLM 12.81 11.32 29.04 17.24 13.67 6.91 16.34 20.25 17.35 41.00 52.74 25.77 0.50 3.00 28.38 30.98  20.46
SnapKV 17.51 17.57 38.89 22.15 17.28 9.13 15.01 21.96 17.94 46.00 61.05 35.97 0.00 4.00 36.92 37.83  24.95
Pyramid 18.17 17.58 39.08 22.05 16.78 8.13 14.74 22.24 17.88 47.50 60.11 37.02 0.50 3.50 36.96 38.73  25.06
Ada-SnapKV 18.64 18.61 39.59 22.51 17.05 9.19 15.28 22.88 18.98 52.50 61.69 36.76 0.00 3.00 36.82 39.63  25.82
Ada-Pyramid 18.35 18.93 39.49 22.57 16.83 8.61 15.05 23.22 18.85 55.50 60.93 37.39 0.50 3.50 36.55 39.79 26.00
B=256h
H2O 18.99 19.18 38.78 21.88 17.33 9.16 16.88 23.29 20.51 48.00 60.36 38.07 3.00 5.50 37.18 37.94  26.00
StreamingLLM 13.59 11.81 29.73 18.59 14.37 6.72 21.06 20.78 21.29 51.50 51.92 26.51 0.50 3.00 28.97 31.09  21.96
SnapKV 19.27 20.61 40.78 22.81 16.83 9.89 16.23 23.17 20.07 53.50 61.75 38.41 0.00 4.00 38.25 40.57  26.63
Pyramid 18.81 19.83 40.71 22.34 17.10 9.08 16.10 22.93 19.50 60.00 61.01 38.65 0.50 5.00 38.23 39.13  26.81
Ada-SnapKV 18.99 21.08 41.18 22.89 17.64 9.52 16.71 23.05 20.48 67.00 61.27 38.74 0.00 3.50 39.60 40.96  27.66
Ada-Pyramid 18.78 20.32 40.50 22.73 17.01 9.37 16.05 23.60 19.93 69.00 61.43 39.07 2.00 5.00 38.40 40.08 27.70
B=512h
H2O 18.61 20.07 39.82 22.08 17.21 10.13 17.62 23.65 21.41 54.50 61.84 38.74 3.00 5.50 39.23 40.08  27.09
StreamingLLM 13.94 13.13 33.06 18.26 14.44 7.41 25.24 21.00 23.78 60.50 52.04 26.31 1.00 3.00 30.10 31.76  23.44
SnapKV 18.45 21.96 42.01 23.25 17.42 9.88 17.68 23.62 21.30 68.00 61.77 39.02 1.00 4.50 40.09 40.79  28.17
Pyramid 18.46 22.85 42.24 23.27 16.75 9.45 17.41 24.62 21.20 70.00 60.61 39.32 3.00 6.50 39.63 40.78 28.51
Ada-SnapKV 18.83 22.39 42.15 23.52 18.27 9.63 17.66 23.99 21.23 70.00 61.72 38.93 2.00 4.50 40.11 41.28 28.51
Ada-Pyramid 18.64 22.86 41.81 23.61 16.67 9.45 17.35 23.75 20.79 70.00 60.66 39.61 3.00 5.50 39.87 40.99  28.41
B=1024h
H2O 17.11 22.34 41.26 22.09 17.47 9.60 18.82 23.94 22.49 61.00 62.33 38.68 3.00 5.50 41.23 41.18  28.00
StreamingLLM 14.78 16.77 37.64 18.77 14.63 7.39 26.43 21.47 24.21 67.00 53.00 25.99 0.50 3.00 31.51 32.31  24.71
SnapKV 18.45 24.18 42.50 23.53 17.32 10.23 19.00 24.26 23.04 69.50 62.22 39.88 3.00 5.50 41.15 41.91 29.10
Pyramid 18.48 24.87 42.11 23.45 16.97 9.84 18.93 24.50 22.77 69.50 61.65 39.73 2.50 5.00 41.07 41.27  28.91
Ada-SnapKV 18.94 23.68 43.27 23.28 17.15 9.89 18.58 23.46 22.65 70.00 62.24 39.83 2.50 5.50 41.68 42.88 29.10
Ada-Pyramid 19.00 23.83 43.36 23.48 17.03 9.32 18.70 24.11 22.61 69.50 61.83 39.75 2.50 6.00 40.85 41.80  28.98
Table 2: Comparison Based on LWM-Text-Chat-1M Among 16 Datasets
A.4 Detailed results for LWM model Among 16 Datasets
The table 2 presents quality scores of different eviction strategies based on the LWM model across 16 datasets. Overall, the results are consistent with those of Mistral, and the adaptive allocation also leads to quality improvements after cache eviction.
A.5 Detailed Information of Datasets
Table 3 provides a comprehensive description of information pertaining to 16 datasets.
Label Task Task Type Eval metric Avg len Language Sample Num
NrtvQA NarrativeQA Single-Doc. QA F1 18,409 EN 200
Qaspe Qasper Single-Doc. QA F1 3,619 EN 200
MF-en MultiFieldQA-en Single-Doc. QA F1 4,559 EN 150
HotpotQA HotpotQA Multi-Doc. QA F1 9,151 EN 200
2WikiMQA 2WikiMultihopQA Multi-Doc. QA F1 4,887 EN 200
Musique MuSiQue Multi-Doc. QA F1 11,214 EN 200
GovReport GovReport Summarization Rouge-L 8,734 EN 200
QMSum QMSum Summarization Rouge-L 10,614 EN 200
MultiNews MultiNews Summarization Rouge-L 2,113 EN 200
TREC TREC Few-shotLearning Accuracy 5,177 EN 200
TriviaQA TriviaQA Few-shotLearning F1 8,209 EN 200
SAMSum SAMSum Few-shotLearning Rouge-L 6,258 EN 200
PCount PassageCount Synthetic Accuracy 11,141 EN 200
PRe PassageRetrieval-en Synthetic Accuracy 9,289 EN 200
Lcc LCC Code Edit Sim 1,235 Python/C#/Java 500
RB-P RepoBench-P Code Edit Sim 4,206 Python/Java 500
Table 3: Information of 16 Datasets
A.6 Detailed Concentration Visualization of Heads
Figure 6 supplements Figure 2 in the main paper by presenting the visualization results across all layers. It can be observed that in all layers, different heads exhibit significant variations in attention concentration. This indicates that the adaptive allocation algorithm has great potential to reduce the upper bound of eviction loss empirically.
A.7 Code
The source code for this paper is intended for public release following the publication. It will be accessible at https://github.com/FFY0/AdaKV once it is made available.