### v2

# Ada-KV: Optimizing KV Cache Eviction by Adaptive Budget Allocation for Efficient LLM Inference

Abstract
Large Language Models have excelled in various fields but encounter efficiency limitations due to the substantial Key-Value (KV) cache required for long-sequence inference. Recent efforts try to evict non-critical cache elements during runtime, thereby reducing cache size within given memory budgets while preserving generation quality. Our reexamination of foundational principles reveals that prevailing methods aim to minimize an upper bound of eviction loss, quantified as the L1 distance between the pre- and post-eviction outputs of multi-head self-attention mechanisms. Moreover, our analysis indicates that the common practices of uniformly assigning budgets across different attention heads during cache eviction hinder their budget utilization, negatively impacting generation quality. In light of these findings, we propose a simple yet effective adaptive budget allocation algorithm. This algorithm not only optimizes the loss upper bound in theory but also reduces the eviction loss in practice by aligning with the intrinsic patterns of self-attention mechanisms. Integrating this algorithm into two advanced methods, we develop Ada-SnapKV and Ada-Pyramid. Extensive evaluations on 16 datasets and the Needle-in-a-Haystack test confirm that they both significantly boost performance across various tasks.
[…]
1 Introduction
Autoregressive Large language models (LLMs) have achieved significant success and are widely utilized across diverse natural language processing applications, including dialogue systems (Yi et al. 2024), document summarization (Laban et al. 2023), and code generation (Gu 2023). The widespread deployments have propelled the development of their capacities to process extended sequences. For instance, GPT-4 supports sequences up to 129K (Achiam et al. 2023), Claude3 up to 200K (Anthropic 2024), and Gemini-Pro-1.5 (Reid et al. 2024) up to 1M tokens. However, transformer-based LLMs pose significant efficiency challenges while processing long-sequence inputs.
Specifically, the inference process of LLMs consists of two stages: prefilling and decoding. During prefilling, LLMs compute and store all Key-Value (KV) cache pairs for the input tokens from the prompt. In the subsequent decoding phase, LLMs autoregressively generate each output token using the most recent token and all previous KV cache pairs. This stage continues until the maximum designated length is reached or it is actively stopped. However, while dealing with the long-sequence inputs, the size of the generated KV cache significantly increases, easily exceeding the model’s parameter size (Sun et al. 2024). This escalation leads to substantial storage challenges during both the prefilling and decoding stages. Additionally, during the decoding phase, each autoregressive step requires extensive KV cache I/O, which incurs significant I/O latency. This latency now exceeds the time required for computation and becomes the primary bottleneck in decoding speed (Tang et al. 2024; Dao et al. 2022). Consequently, the extensive KV cache associated with long sequences poses significant efficiency challenges for the inference process.
In response to the challenges posed by large KV cache size, a range of KV cache eviction methods have been developed, as referenced in recent literature (Zhang et al. 2024b; Xiao et al. 2023; Ge et al. 2023; Yang et al. 2024; Zhang et al. 2024a; Li et al. 2024). As shown in Figure 1, these methods discard most non-critical cache pairs after the prefilling stage of each layer in LLMs. Therefore, by tailoring the cache sizes to specified storage budgets, they not only reduce the storage burden but also boost the decoding speed while maintaining the post-eviction generation quality. These eviction methods are designed with plug-and-play capabilities, allowing for straightforward integration into any LLM setup without the need for fine-tuning. They typically utilize various strategies to selectively evict the majority of KV cache pairs. The latest leading methods (Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a), employ a Top-K based selection scheme based on attention weights. These approaches effectively distinguish between critical and non-critical cache pairs, deciding which to retain and which to evict. Despite these advancements, the challenge of minimizing quality loss while employing these eviction methods remains an unresolved issue in the field.
Figure 1: Cache Eviction Framework. Cache Eviction removes a large number of non-essential tokens, marked by dashed borders, immediately after the prefilling stage of each layer. This process substantially reduces storage requirements and accelerates decoding, while preserving the quality of post-eviction generations.
Our study begins by reexamining the underlying principles of eviction methods from a theoretical perspective. We reveal that Top-K based methods aim to minimize an upper bound of eviction loss, quantified as the L1 distance between the pre- and post-eviction outputs of self-attention mechanisms. Moreover, We find the common practice of uniformly assigning budgets across different attention heads leads to a misallocation of budgets, thus limiting budget utilization. Based on these insights, we propose a simple yet effective adaptive allocation algorithm. As an illustrative example shown in Figure 2, it effectively improves budget utilization by dynamically allocating the budget across different attention heads within the same layer based on their concentration degrees, thereby enhancing the quality of post-eviction generation. We substantiate the advantages of adaptive allocation both theoretically and empirically. Theoretically, as detailed in Section 3.4, the adaptive allocation reduces the upper bound of eviction loss compared to the current uniform allocation. Empirically, as detailed in Section 3.5, the alignment of the adaptive allocation with inherent varied degrees in attention concentration among the heads within self-attention mechanisms effectively reduces eviction loss in practice.
By integrating this adaptive budget allocation algorithm into two advanced methods (Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a), we develop two adaptive eviction methods: Ada-SnapKV and Ada-Pyramid, respectively. Extensive evaluations across 16 datasets, spanning various tasks in LongBench (Bai et al. 2023), demonstrate that both Ada-SnapKV and Ada-Pyramid significantly improve the generation quality. Additionally, the Needle-in-a-Haystack test, a widely used evaluation method, further confirms that adaptive allocation enhances in-context retrieval ability. The main contributions are summarized as follows:
  • Reexamination of Cache Eviction Principle: We reveal the principles of cache eviction methods that aim to minimize an upper bound of eviction loss, quantified as the L1 distance between the pre- and post-eviction attention outputs.
  • Identifying and Addressing Limitations in Budget Allocations: We pinpoint the inefficiency of existing uniform budget allocation across heads. To address this, we introduce an adaptive allocation algorithm that not only optimizes the upper bound in theory but also reduces the eviction loss in practice.
  • Developments of Adaptive Eviction Methods: By integrating adaptive allocation, we develop two novel methods: Ada-SnapKV and Ada-Pyramid. These methods both significantly improve the post-eviction generation quality in the extensive evaluations.
  • Innovative Direction to Enhance Cache Eviction: Our research pioneers a new direction by adaptively allocating budgets based on unique characteristics among heads to optimize cache eviction. This innovation, guided by our theoretical analysis of the upper bounds of eviction loss, seeks significant further development in this field.
(a) Uniform Allocation
(b) Adaptive Allocation
Figure 2: An Illustrative Example.This example features a five KV cache pairs with associated attention weights across three heads. Adaptive allocation, through reallocating budgets from the concentrated Head2 to the dispersed Head1, increases the aggregated weights of retained KV cache pairs from 2.33 to 2.46, which correlates to a reduced eviction loss in Section 3.4 and 3.5.
2 Related Works
In the context of long-sequence inference, the vast scale of the KV cache pairs leads to a memory-bound situation, causing significant storage and I/O latency costs  (Wang and Chen 2023). Some studies employ efficient memory management strategies that reduce I/O time without altering the size of the KV Cache, such as Page Attention (Kwon et al. 2023) and Flash Attention (Dao et al. 2022). These studies are orthogonal to our work and we incorporate the Flash Attention technique in our implementation to achieve efficient computation. Some works, called KV cache quantization, reduce the size of cache pairs by lowering the precision of individual entries. However, its compression ratio is inherently limited—ideally, quantization from 16-bits to 1-bit achieves a theoretical limit of 6.25% compression but results in catastrophic quality degradation. However, cache eviction methods could compress the KV Cache to any given small budget, like less than 1% of the original size while still maintaining certain generation quality. Moreover, the techniques of quantization and eviction are orthogonal, suggesting potential for further combined optimization in the future. Another work (Tang et al. 2024) attempts to reduce I/O overhead by only recalling KV cache pairs relevant to the current query for computation. However, it is constrained by substantial memory overheads, making deployment on GPUs with limited storage capacities challenging.
Recently, KV cache eviction methods have gained attention and rapid development due to their flexible compression ratios given any storage budgets and the advantage of being plug-and-play without the need for fine-tuning. Earlier StreamingLLM (Xiao et al. 2023) simply maintains the cache of 4 initial and the recent tokes, discarding all others to adapt to long sequence inference. FastGen (Ge et al. 2023) searches and combines multiple strategies, like maintaining the cache of special tokens, punctuation tokens, and recent tokens, based on the characteristics of attention heads. H2O (2024b) has developed an eviction algorithm that utilizes query states of all tokens to identify important KV pairs based on the Heavy Hitter method. The most recent SnapKV (Li et al. 2024) identifies important KV cache pairs using query states of several tokens within a recent window, evicting the less important ones. This method effectively mitigates the quality degradation in cache eviction. The Pyramid (Yang et al. 2024; Zhang et al. 2024a) further adjusts the budget allocation across different layers in SnapKV, improving the generation quality in small-budget scenarios. However, to our best knowledge, current eviction methods have never tried to adaptively distribute the total budget across different heads. Building on theoretical analysis and observations of inherent attention patterns in LLM heads, we identify and demonstrate the necessity for adaptive budget allocation in cache eviction.
3 Framework
3.1 Overview
In this section, we reexamine how the existing cache eviction strategies retain essential information in the past KV cache from a theoretical perspective. Inspired by theoretical findings, we propose a simple yet effective algorithm for adaptive budget allocation, which is proven to be better than the previous uniform allocation method in cache eviction procedures, both theoretically and practically. Upon further integrating this algorithm into two leading existing methods, we have developed two adaptive cache eviction techniques: Ada-SnapKV and Ada-Pyramid. The key findings and insights are as below:
  • How do the eviction strategies maintain generation quality despite discarding substantial KV cache? Theorem 2 formalizes an upper bound of the eviction loss by using the L1 distance between the pre- and post-eviction outputs and Theorem 3 demonstrates that Top-K based eviction strategies aim to minimize this upper bound.
  • Why is adaptive budget allocation better from the theoretical perspective? Theorem 4 demonstrates that the upper bound of eviction loss under adaptive allocation maintains at or below that under previous uniform allocation in the cache eviction procedure. Thus it implies that the adaptive allocation is better in theory.
  • Why is adaptive budget allocation better from the empirical perspective? As shown in Figure 3 and 4, we identify that the inherent disparities in attention concentration across different heads reliably translate theoretical advantages into practical results, leading to consistent reductions of eviction loss in practice.
3.2 Preliminary
We begin by providing a formal description of the computational processes involving Multi-head Attention and KV cache in a single-layer of LLMs to alleviate the burden of notation. LLMs are characterized by their autoregressive generation mode, where each step involves using the last token to predict the next token. Define $X\in\mathbb{R}^{n\times d}$ as the embedding matrix of all tokens in sequence, and $x\in\mathbb{R}^{1\times d}$ as the last token used as input at the current timestep. To clarify the subsequent theoretical exposition, we adopt the notation system from  (Liu et al. 2023) under the assumption of $h$ attention heads in one layer. The transformation matrices for each head $i\in[h]$, $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\,\in\mathbb{R}^{d\times d_{h}}$, map token embeddings to their respective Query, Key, and Value and the final output matrix $W_{i}^{O}\,\in\mathbb{R}^{d_{h}\times d}$ transform the intermediate result to the output hidden states. At each timestep, the previous KV cache pairs on head $i$ have been initialized as:
$K_{i}=XW_{i}^{K},V_{i}=XW_{i}^{V}$ (1)
Then, the input token $x$ is mapped to its corresponding query, key, value for each head, and the previous KV cache is updated accordingly:
$q_{i}=xW_{i}^{Q},k_{i}=xW_{i}^{K},v_{i}=xW_{i}^{V}$ (2)
$K_{i}=[K_{i}:k_{i}],V_{i}=[V_{i}:v_{i}]$ (3)
Finally the final output $o\in\mathbb{R}^{1\times d}$ is computed as follow:
$o=\sum^{i\in[h]}A_{i}V_{i}W_{i}^{O}$ (4)
where $A_{i}\in\mathbb{R}^{1\times n}$ is the attention weight calculated by [^1] The scaling factor $\sqrt{d_{h}}$ is omitted for simplification.:
$A_{i}=\text{softmax}(q_{i}K_{i}^{T})$ (5)
3.3 Reexamining the Principle of Cache Eviction
Cache eviction is dedicated to reducing the size of KV cache to fit within a constrained budget by evicting non-critical cache pairs strategically. Eviction masks $\left\{\mathcal{M}_{i}\in\mathbb{R}^{1\times n}\right\}$ can be employed to simulate the post-eviction output $o^{\prime}$ of self-attention mechanism during the generation process:
$o^{\prime}=\sum^{i\in[h]}A_{i}^{\prime}V_{i}W_{i}^{O}\>\text{where}\>A_{i}^{\prime}=\text{softmax}(q_{i}K_{i}^{T}+\mathcal{M}_{i})$ (6)
where each element $\mathcal{M}_{i}^{j}$  [^2] Given the first dimension of $\mathcal{M}_{i}$ is 1, $M_{i}^{j}$ is used to simplify the notation for $\mathcal{M}_{i}(1,j)$. Similarly, $A_{i}^{j}$ is in the same manner. in mask $\mathcal{M}_{i}$ indicates that whether the $j\text{th}\in[n]$ KV cache pair has been evicted in $K_{i},V_{i}\in\mathbb{R}^{n\times d_{h}}$ on each head $i$:
$\mathcal{M}_{i}^{j}$ =$=\begin{cases}0&\text{if the $j$th cache pair on head $i$ is retained }\\ -\infty&\text{otherwise the $j$th cache pair on head $i$ is evicted}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
Thus, the budget $B_{i}$ for head $i$ corresponds to the number of zero elements in $\mathcal{M}_{i}$. Theorem 1 further simplifies the output $o^{\prime}$ by eliminating the softmax function. A detailed proof is provided in Appendix A.1.
Theorem 1.
The post-eviction output $o^{\prime}$ can rewrite as:
$o^{\prime}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (7)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
[…]
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
The reduction in generation quality during cache eviction stems from changes in the attention output. Thus, we quantify the eviction loss as the L1 distance between the pre- and post-eviction outputs of self-attention mechanisms:
Eviction $\text{Eviction Loss}=||o-o^{\prime}||_{1}$ (8)
Utilizing the row norm of matrix, we derive an upper bound $D$ for the Eviction Loss in Theorem 2. For a detailed proof, refer to Appendix A.2.
Theorem 2.
The eviction loss caused by cache eviction can be bounded by $D$ as follows:
Eviction $\text{Eviction Loss}\leq D=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (9)
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
where $C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{\infty}\right\}$ is the max value in the row norms of Matrices $\left\{V_{i}W_{i}^{O}\right\}$ among all heads.
Cache eviction strategies typically presuppose the stability of critical cache pairs (Zhang et al. 2024b; Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a), thereby facilitating cache eviction during the autoregressive generation process. Consequently, the advanced strategies leverage query states from several recent tokens to calculate attention weights with all past cache pairs, which, in conjunction with Top-K selections, approximate the identification of cache pairs critical in subsequent generations. For simplicity, we assume that the eviction procedure relies solely on a single query state $q_{i}$ associated with the last token $x$ for critical cache detection. As shown in Algorithm 1, the core idea of current leading eviction strategies based on Top-K selections is only to retain cache pairs corresponding to $B_{i}$ highest weights $A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})$, while evicting those considered non-essential. Obviously, given the budget allocation $\left\{B_{i}\right\}$ the Top-K based eviction strategies maximize the aggregated weights as follows:
[…]
given budget $\text{given budget allocation}\>\left\{B_{i}\right\}\>\text{s.t.}\>\sum^{i\in[h]}B_{i}=B$
Therefore, we establish Theorem 3, which demonstrates that the principle of Top-K based eviction aims to minimize an upper bound $D$.
Algorithm 1 Eviction Based on Top-K Selection
Input: Allocation $\left\{B_{i}\right\}$,KV Cache $\left\{K_{i},V_{i}\right\}$, Weights $\left\{A_{i}\right\}$
Output: Compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
1:  for $i\leftarrow 1$ to $h$ do
2:     initialize empty cache $\hat{K}_{i},\hat{V}_{i}$ for head $i$
3:     for $j\leftarrow 1$ to $n$ do
4:        if $A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})$ then
5:           retain and append cache pair $K_{i}^{j},V_{i}^{j}$ to $\hat{K}_{i},\hat{V}_{i}$
6:        else
7:           evict cache pair $K_{i}^{j},V_{i}^{j}$
8:        end if
9:     end for
10:  end for
11:  return  compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
Theorem 3.
Given a budget allocation result $\left\{B_{i}\right\}$, the principle of eviction strategies based on Top-K selections is to minimizes the upper bound $D$ of eviction loss.
[…]
given budget $\>\text{given budget allocation}\>\left\{B_{i}\right\}\>\text{s.t.}\>\sum^{i\in[h]}B_{i}=B$
3.4 Adaptive vs. Uniform:Theoretical Perspective
In existing studies on cache eviction, the budget $B$ is uniformly assigned across all heads within a layer: specifically, ${B_{i}=B/h}$. Consequently, the upper bound of eviction loss $D$ is modified to $D^{\prime}$ under uniform allocation:
$D^{\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}$)}}A_{i}^{j}$ (12)
given uniform budget $\text{given uniform budget allocation}\left\{B_{i}=B/h\right\}$
In contrast, we suggest adaptive allocation of the overall budget among heads and introduce a simple yet effective budget allocation algorithm that dynamically distributes the total budget $B$ based on the attention weights $A_{i}$ across heads, with the allocated results denoted as $\left\{B^{*}_{1},B^{*}_{2},...,B^{*}_{h}\right\}$ subject to $\sum^{i\in[h]}B^{*}_{i}=B$. As shown in Algorithm 2, it firstly selects the $B$ largest attention weights from all heads within one layer. Based on the times each head is selected during the above procedure, different budgets ${B^{*}_{i}}$ are allocated to each head. Under this adaptive allocation, the upper bound of eviction loss $D$ is modified to $D^{\prime\prime}$:
$D^{\prime\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}^{*}$)}}A_{i}^{j}$ (13)
given adaptive budget $\text{given adaptive budget allocation}\left\{B_{i}=B_{i}^{*}\right\}$
Theorem 4.
The upper bound $D^{\prime\prime}$ of eviction loss with adaptive budget allocation consistently remains at or below the upper bound $D^{\prime}$ associated with uniform allocation.
$D^{\prime\prime}\leq D^{\prime}$ (14)
Algorithm 2 Adaptive Budget Allocation
Input: Total Budget $B$, Attention Weights in $h$ heads $\left\{A_{i}\in\mathbb{R}^{1\times n}\right\}$;
Output: Allocated Budgets of $h$ heads $\left\{B^{*}_{i}\right\}$;
1:  Concatenate across heads $A$ = Cat($\left\{A_{i}\right\}$,dim=1)
[…]
3:  Identify top indices $T$ = Top-K($A,B$).indices
4:  Select the corresponding head indicator $I^{*}=I[T]$
5:  Count frequencies of each $i$ in $I^{*}$ to determine $\left\{B^{*}_{i}\right\}$
6:  return  Allocated Results of $h$ heads $\left\{B^{*}_{i}\right\}$
According to Theorem 4, our adaptive allocation algorithm achieves equal or smaller eviction loss upper bound compared to the previous uniform allocation, thereby offering theoretical advantages. The detailed proof can be found in Appendix A.3.
Figure 3: Disparities in Attention Concentration Across Heads. (Mistral-instruct-v0.2 on the first sample of the Qasper, a single-doc QA dataset in LongBench) We aggregate different proportions of the top attention weights, $\sum^{\text{retained}j}A_{i}^{j}$, to analyze attention concentration in different head $i$. Most concentrated heads require a small cache proportion, e.g., $5\%$, to aggregate weights close to 1, whereas other dispersed heads need significantly larger proportions, such as $50\%$.
3.5 Adaptive vs. Uniform: Empirical Perspective
According to Theorem 4, the upper bound of eviction loss of adaptive allocation is equal to or less than that of the prevailing uniform allocation. We further demonstrate different attention heads within each layer of LLMs exhibit significant disparities in the degrees of attention concentration, which effectively translates this theoretical advantage of a reduced upper bound into practical reductions in eviction loss.
For visualization in Figure 3 (a), most concentrated heads in Layer 8 like head 1 require only 1% of the original cache budget to effectively retain the aggregated weights $\sum^{\text{retained}j}A_{i}^{j}$ of 0.95. Conversely, other heads like head 18 require nearly 50% proportion to near 0.95. This characteristic is closely related to the upper bound $D$ of eviction loss, as detailed in Theorem 3. Under such circumstances, the previous uniform allocation encounters a budget utilization dilemma, as an illustrative example in Figure 2: it must either endure substantial eviction losses in dispersed heads or waste excessive and unwarranted budgets on heads with concentrated attention. This significantly undermines the trade-off performance between the total budget and generation quality in the cache eviction. In contrast, the adaptive allocation algorithm assigns large budgets to dispersed heads, while controlling the budget sizes for other heads, effectively maintaining the overall size and mitigating the decline in generation quality. In Figure 4, we show the averaged relative eviction loss in all layers, $\frac{\lVert o^{\prime}-o\rVert_{1}}{\lVert o\rVert_{1}}$, with uniform and adaptive allocations separately. Results show that the adaptive allocation consistently reduces eviction losses across all 200 samples in the Qasper dataset. This indicates that adaptively allocating budgets among different heads enhances the budget utilization significantly, thereby reducing the output eviction loss.
Algorithm 3 Ada-SnapKV/Ada-Pyramid in One Layer
Input: Overall budget $B$, Past cache $\left\{K_{i},V_{i}\right\}$, Tokens in the recent window $X^{rec}\in\mathbb{R}^{win*d}$
Output: Compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
1:  for $i\leftarrow 1$ to $h$ do
2:     $Q_{i}^{rec}=X^{rec}W_{i}^{Q}$
3:     $\bar{A_{i}}=softmax(Q_{i}^{rec}K_{i}^{T})$
4:     $\bar{A_{i}}=\bar{A_{i}}.maxpooling(dim=1).mean(dim=0)$
5:  end for
6:  get $\left\{B_{i}^{*}\right\}$ by invoking Algorithm 2($B,\left\{\bar{A}_{i}\right\}$)
7:  $\left\{B_{i}^{*}\right\}=\alpha\times\left\{B_{i}^{*}\right\}+(1-\alpha)\times\frac{B}{h}$
8:  $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$ = Algorithm 1($\left\{B_{i}^{*}\right\},\left\{K_{i},V_{i}\right\},\left\{\bar{A}_{i}\right\}$)
9:  return  compressed cache $\left\{\hat{K}_{i},\hat{V}_{i}\right\}$
(a) $B$ = 5% original size
(b) $B$ = 10 % original size
Figure 4: Comparison of Eviction Losses. (Using Mistral-7B-instruct-v0.2 on 200 samples from the Qasper Dataset) Cache eviction is implemented under Uniform and Adaptive allocation strategies respectively, compressing the original cache size to 5% and 10%. The adaptive version consistently yields a lower average relative eviction loss $\frac{\lVert o-o^{\prime}\rVert_{1}}{\lVert o\rVert_{1}}$ across all layers on different samples compared to the prevailing uniform allocation.
3.6 Implementation
The current two leading eviction strategies, SnapKV and Pyramid, both utilize the several tokens $X^{rec}\in\mathbb{R}^{win*d}$ from a recent window (typically window size is 32) to identify and evict the less important cache pairs. SnapKV excels at scenarios with large budgets, while Pyramid is more effective in smaller budgets. The key distinction lies in how Pyramid and SnapKV allocate budgets within different layers in LLMs. Pyramid suggests that information aggregation among layers takes a pyramidal form, thereby allocating a larger budget to shallower layers and progressively reducing it in deeper layers through pre-set hyper-parameters. In contrast, SnapKV distributes the budget uniformly across all layers. Nonetheless, same as other eviction methods, they both allocate the budget uniformly among all heads within a single layer. Thus, for a specific layer with a overall budget of $B$, their eviction algorithms are the same.
We incorporate our adaptive allocation algorithm into both of them, resulting in the creation of two novel adaptive strategies: Ada-SnapKV and Ada-Pyramid, respectively. As shown in Algorithm $3$Algorithm 3 describes the process serially for simplicity, but eviction operations can be parallelized easily in practice., the adaptive budget allocation can be seamlessly integrated into any eviction strategy by invoking the adaptive allocation Algorithm 2 before eviction process in each layer. Both of Ada-SnapKV and Ada-Pyramid utilize a max pooling layer to facilitate cache selection following the previous SnapKV (Li et al. 2024) and Pyramid (Zhang et al. 2024a). Besides, a hyper-parameter $\alpha$ in line 6, set by default to 0.5, prevents assigning tiny budgets to highly concentrated heads, thereby enhancing fault tolerance of critical cache detection. Additionally, we incorporate the Flash Attention (Kwon et al. 2023) technique to support efficient computation of variable-length head caches, with the computational efficiency evaluations provided in Appendix A.4.
Single-Doc. QA Multi-Doc. QA Summarization Few-shotLearning Synthetic Code
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Ave. Score
Full Cache 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 55.33 52.87   42.51
B=128h
H2O 21.19 21.66 38.60 30.63 20.65 12.19 20.65 22.42 21.81 39.00 82.52 40.68 2.98 79.56 49.13 46.76  34.40
StreamingLLM 16.61 14.74 31.40 28.05 21.36 12.08 18.44 18.91 19.26 43.50 74.22 29.00 2.75 31.65 41.27 38.84  27.63
SnapKV 19.17 21.40 42.93 36.76 22.44 15.86 19.16 21.84 21.55 47.50 84.15 40.24 2.30 68.26 50.69 47.13  35.09
Pyramid 20.16 21.77 43.55 36.78 23.12 14.39 19.53 22.03 21.47 51.00 84.62 40.24 2.79 70.77 50.57 46.53  35.58
Ada-SnapKV 20.63 22.58 45.68 37.90 23.49 16.55 19.99 22.28 21.55 59.50 85.00 40.62 3.09 69.36 50.98 48.17  36.71
Ada-Pyramid 20.50 21.71 45.61 36.81 23.57 15.84 19.75 22.13 22.00 60.50 84.04 40.51 3.21 73.60 51.24 48.02 36.81
B=256h
H2O 21.54 22.92 42.56 31.07 22.53 13.76 22.52 22.40 23.09 40.50 84.20 40.77 3.41 86.10 50.98 48.17  36.03
StreamingLLM 17.93 16.01 33.36 30.71 21.30 10.08 20.66 19.47 22.89 53.50 73.59 29.22 3.00 27.77 42.30 39.87  28.85
SnapKV 22.37 23.74 48.13 38.56 22.43 15.66 21.91 23.13 23.15 61.50 85.45 41.42 3.09 84.54 53.22 50.24  38.66
Pyramid 20.09 24.00 47.33 38.24 22.48 16.02 21.40 22.45 22.63 63.00 84.93 40.98 3.40 82.48 52.78 49.36  38.22
Ada-SnapKV 22.55 25.78 48.33 40.30 24.24 16.64 21.63 23.03 23.19 67.00 85.78 41.53 3.47 87.07 53.86 51.13 39.72
Ada-Pyramid 22.64 24.64 47.40 40.25 23.62 16.83 21.82 23.34 22.70 66.50 84.99 41.34 2.78 86.90 53.17 49.52  39.28
B=512h
H2O 21.72 26.03 44.81 32.33 23.16 14.86 23.65 22.84 24.70 42.00 85.22 41.57 3.40 86.45 53.04 49.68  37.22
StreamingLLM 18.76 17.17 37.09 30.21 21.64 9.93 24.44 20.00 25.57 62.00 72.36 29.95 2.48 18.17 43.70 40.13  29.60
SnapKV 24.60 27.81 48.98 39.46 25.25 16.98 23.70 22.96 24.37 67.00 85.88 41.26 2.78 86.56 54.81 51.71  40.26
Pyramid 23.23 27.94 48.87 40.50 24.36 16.74 23.22 23.16 24.37 67.00 85.73 41.74 3.16 85.67 54.16 50.34  40.01
Ada-SnapKV 23.39 28.72 48.96 40.60 25.20 17.25 23.15 23.48 24.41 68.00 86.39 41.69 2.73 88.92 54.69 51.51 40.57
Ada-Pyramid 24.03 28.98 48.39 39.25 24.50 18.38 23.13 23.90 24.30 68.00 85.89 41.89 2.98 87.71 54.46 51.39  40.45
B=1024h
H2O 23.90 28.62 46.46 37.03 24.74 15.04 25.30 23.11 25.92 46.00 85.93 41.80 3.24 86.57 54.46 51.01  38.70
StreamingLLM 19.42 21.69 41.75 32.40 22.18 11.18 27.13 21.09 26.59 67.00 71.79 30.11 2.88 16.57 44.82 39.76  31.02
SnapKV 25.47 29.57 49.33 40.90 25.53 19.01 25.94 23.89 26.21 69.50 86.48 42.10 2.98 88.56 55.57 51.92  41.44
Pyramid 24.21 29.86 48.93 40.75 25.05 18.77 25.73 24.06 25.65 68.50 86.31 42.25 2.97 87.17 54.75 52.10  41.07
Ada-SnapKV 24.79 31.94 48.45 40.73 26.22 19.11 25.61 23.92 26.03 70.00 86.32 42.35 2.91 88.31 55.44 52.55 41.54
Ada-Pyramid 25.09 30.94 48.18 40.00 26.52 19.10 24.93 23.71 25.86 70.00 86.34 42.64 2.56 86.92 54.93 51.90  41.23
Table 1: Comparison Based on Mistral-7B-Instruct-v0.2 Among 16 Datasets
(a) Mistral-7B-Instruct-v0.2
(b) LWM-Text-Chat-1M
Figure 5: Average Score Among 16 Datasets
(a) SnapKV
(b) Ada-SnapKV
(c) Pyramid
(d) Ada-Pyramid
Figure 6: Needle-in-a-Haystack Test. This test inserts a critical sentence (the ’needle’) within a document (the ’haystack’) with extensive context, then evaluates a model’s ability to retrieve the needle from the document. The x-axis indicates the context length of the document, and the y-axis shows the insertion depth of the needle. The Average Score is determined by averaging the aggregated scores at various context lengths. Higher scores indicate an improved capacity of the model for contextual retrieval.
4 Experiments
4.1 Settings
Datasets
Firstly, we carry out an comprehensive evaluation using 16 datasets, covering domains of single-document QA (Kočiskỳ et al. 2018; Dasigi et al. 2021), multi-document QA (Yang et al. 2018; Ho et al. 2020; Trivedi et al. 2022), summarization (Huang et al. 2021; Zhong et al. 2021; Fabbri et al. 2019), few-shot learning (Joshi et al. 2017; Gliwa et al. 2019; Li and Roth 2002), Synthetic (Bai et al. 2023), and code generation (Guo et al. 2023; Liu, Xu, and McAuley 2023), within LongBench (Bai et al. 2023), a benchmark for evaluating multi-task performance with long-sequence inputs. These datasets feature varying average input lengths from 1,235 to 18,409 tokens, necessitating substantial KV cache size during generation, thereby rendering them suitable for evaluating KV cache eviction strategies. Each dataset is assessed using LongBench-recommended metrics, with quality scores up to 100. Detailed dataset information is provided in Appendix A.6. We also utilize the widely-used ’Needle-in-a-Haystack’ test, where key information is randomly inserted into long texts to create prompts. This test evaluates whether LLMs can extract this key information from extensive texts, specifically examining the impact of proposed adaptive allocation on the models’ fundamental long context retrieval abilities.
Baselines
We conduct a comparative analysis between our newly proposed methods, Ada-SnapKV and Ada-Pyramid, and the previously established SnapKV and Pyramid, acknowledged as two leading approaches under varying scenarios. Additionally, we assess these methods against other earlier eviction strategies, including StreamingLLM and H2O. StreamingLLM is designed to handle long-sequence inputs by retaining caches of several initial tokens and others in a recent window. And H2O compresses cache size by employing attention weight-based detection of Heavy Hitters based on all queries and key states. In all experiments, the hyper-parameter $\alpha$ in Algorithm 2 is set to 0.5. Both Ada-SnapKV and Ada-Pyramid, as well as SnapKV and Pyramid, utilize the same configuration settings as described in (Li et al. 2024), ensuring comparability with a recent window size of 32 and a maximum pooling kernel size of 7. Parameters for StreamingLLM and H2O conform to the default settings reported in the literature (Zhang et al. 2024b; Xiao et al. 2023).
Base Models
In the experiments, we employ two open-source base models: Mistral-7B-instruct-v0.2 (Jiang et al. 2023) and LWM-Text-Chat-1M (Liu et al. 2024). The Mistral 7B model features a context length of 32K and has been adopted as the primary model in related studies (Li et al. 2024; Zhang et al. 2024a) due to its moderate parameter size and remarkable capability for long text tasks. Meanwhile, LWM 7B model stands as the state of the art with its 1M context length, facilitating performance evaluations for the Needle-in-a-Haystack test under extreme context lengths.
4.2 Evaluations Among 16 Datasets
We assess all eviction strategies using cache budget $B\in\left\{128\times h,256\times h,512\times h,1024\times h\right\}$ for each layer. Detailed results for each dataset on the Mistral model are provided in Table 1, while other results for the LWM model are placed in Appendix A.5 due to space constraints. To demonstrate the efficacy of adaptive allocation, we take a budget $B=128h$ as an example presented in Table 1. After integrating the adaptive allocation algorithm, Ada-SnapKV enhances the quality scores in 15 out of 16 datasets compared to the original SnapKV, increasing the average score from 35.09 to 36.71. Similarly, Ada-Pyramid surpasses the original Pyramid in 14 of 16 datasets, boosting the average score from 35.58 to 36.81.
Figure 5 summarizes average scores of all strategies based on Mistral and LWM across 16 datasets. Overall, SnapKV and Pyramid, as the current two leading methods, exhibit closely matched performance, surpassing previous approaches such as H2O and StreamingLLM. Furthermore, our Ada-SnapKV and Ada-Pyramid strategies consistently improve the generated quality under varying budgets, especially in small budgets. The two adaptive eviction strategies alternatingly lead and surpass previous versions to become the new state-of-the-art methods. This consistent improvement validates the necessity and effectiveness of adaptive budget allocation, as previously demonstrated in both theoretical derivations and empirical findings.
4.3 Evaluations on Needle-in-a-Haystack Test
As shown in Figure 6, we employ a Needle-in-a-Haystack test to demonstrate how adaptive budget allocation can enhance long-context retrieval capabilities. All configurations maintain a recent window size of 32 and a pooling kernel size of 7 which is consistent with former experiments, where the maximum inference length is limited to 37K in the full cache case on A800-80G. With a cache budget of $B=128h$, the Ada-SnapKV and SnapKV extend the max length up to 429K, while the Ada-Pyramid and Pyramid extend to 365K. Notably, Ada-SnapKV and Ada-Pyramid both effectively improve long-text retrieval capabilities compared to previous SnapKV and Pyramid. In particular, Ada-SnapKV and Ada-Pyramid achieve near-lossless retrieval within the original 37K length, a feat not replicated by the standard SnapKV and Pyramid. In terms of average score, Ada-SnapKV improves from 94.84 to 95.99, while Ada-Pyramid increases from 96.02 to 96.69. Further details on decoding latency and maximum memory consumption during inference within varied context lengths are available in Appendix A.4.
5 Conclusion
In this study, we reexamine prevailing cache eviction strategies for efficient inference in long sequences, revealing their goal to minimize an upper bound of eviction loss, quantified as the L1 distance between pre- and post-eviction outputs. By introducing an adaptive budget allocation among various attention heads, we theoretically reduce this upper bound compared to previous methods. Our empirical findings suggest that this adaptive approach significantly benefits from the varied concentration degrees inherent in multiple heads within the self-attention mechanism. The development of two novel adaptive eviction methods, Ada-SnapKV and Ada-Pyramid, incorporating this adaptive allocation, demonstrates significant improvements in comprehensive evaluations. Furthermore, this research highlights the substantial potential for advancing cache eviction strategies through our theoretical framework and adaptive budget allocation, specifically designed to exploit the unique characteristics of different attention heads in LLMs.
Appendix A Appendix
A.1 Proof of Theorem 1
Theorem.
The post-eviction output $o^{\prime}$ can rewrite as:
$o^{\prime}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (15)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
[…]
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
Proof.
Consider the softmax function as:
$softmax(x)=\frac{exp(x^{j})}{\sum^{j}exp(x^{j})}$ (16)
Thus, the attention weight after eviction procedure is:
$A^{\prime}_{i}=softmax(s_{i}+\mathcal{M}_{i})\>where\>s_{i}=q_{i}K_{i}^{T}$ (17)
$A^{\prime}_{i}$ $=\frac{exp(s_{i}^{j}+\mathcal{M}_{i}^{j})}{\sum^{j}exp(s_{i}^{j}+\mathcal{M}_{i}^{j})}=\frac{exp(s^{j}_{i})\odot N_{i}^{j}}{\sum^{j}exp(s^{j}_{i})\odot\mathcal{N}_{i}^{j}}$ (18)
$=\frac{exp(s^{j}_{i})\odot N_{i}^{j}}{\sum^{j}exp(s^{j}_{i})}\frac{\sum^{j}exp(s^{j}_{i})}{\sum^{j}exp(s^{j}_{i})\odot\mathcal{N}_{i}^{j}}$ (19)
$=\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}$ (20)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
Thus:
$o^{\prime}=\sum^{i\in[h]}A_{i}^{\prime}V_{i}W_{i}^{O}=\sum^{i\in[h]}\frac{A_{i}\odot\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}}V_{i}W_{i}^{O}$ (22)
$\text{where}\>\mathcal{N}_{i}^{j}=\begin{cases}1&\text{if $K_{i}^{j}\>\text{and}\>V_{i}^{j}$ are retained}\\ 0&\text{otherwise, evict $K_{i}^{j}\text{and}V_{i}^{j}$}\\ \end{cases}$
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
A.2 Proof of Theorem 2
Theorem.
The eviction loss caused by cache eviction can be bounded by $D$ as follows:
Eviction $\text{Eviction Loss}\leq D=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (23)
$\text{given budget allocation}\left\{B_{i}\right\}\>s.t.\>{\sum^{i\in[h]}B_{i}=B}$
where $C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{1}\right\}$ is the max value in the L1-norm of Matrices $\left\{V_{i}W_{i}^{O}\right\}$ among all head.
Proof.
By calculating the L1 distance between their outputs, we can obtain
$||o^{\prime}-o||_{1}=||\sum^{i\in[h]}(\textbf{1}-\frac{\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}V_{i}W_{i}^{O}||_{1}$ (24)
$\leq\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{||A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}V_{i}W_{i}^{O}||_{1}$ (25)
$\leq\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{\lVert A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}\rVert_{1}\>\lVert V_{i}W_{i}^{O}\rVert_{\infty}$ (26)
$\leq C\sum^{i\in[h]}||(\textbf{1}-\frac{\mathcal{N}_{i}}{\lVert A_{i}\odot\mathcal{N}_{i}||_{1}})\odot A_{i}\rVert_{1}$ (27)
$where\>C=Max\left\{\lVert V_{i}W_{i}^{O}\rVert_{\infty}\right\}$
By expanding $A_{i}$, we can further simplify the expression.
$||o^{\prime}-o||_{1}\leq C\sum^{i\in[h]}\sum^{\text{retained}j}(\frac{A_{i}^{j}}{\sum^{\text{retained}j}A_{i}^{j}}-A_{i}^{j})+\sum^{\text{evicted}j}A_{i}^{j}$ (28)
$=C\sum^{i\in[h]}\sum^{\text{retained}j}(\frac{A_{i}^{j}}{\sum^{\text{retained}j}A_{i}^{j}})-\sum^{\text{retained}j}A_{i}^{j}+\sum^{\text{evicted}j}A_{i}^{j}$ (29)
$=C\sum^{i\in[h]}(2-2\sum^{\text{retained}j}A_{i}^{j})$ (30)
$=2hC-2C\sum^{i\in[h]}\sum^{\text{retained}j}A_{i}^{j}$ (31)
A.3 Proof of Theorem 4
Theorem.
The upper bound $D^{\prime\prime}$ of eviction loss with adaptive budget allocation consistently remains at or below the upper bound $D^{\prime}$ associated with uniform allocation.
$D^{\prime\prime}\leq D^{\prime}$ (32)
Proof.
$D^{\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}$)}}A_{i}^{j}$ (33)
given uniform $\text{given uniform allocation}\left\{B_{i}=B/h\right\}$
$D^{\prime\prime}=2hC-2C\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{ Top-K($A_{i}$,$B_{i}^{*}$)}}A_{i}^{j}$ (34)
given adaptive $\text{given adaptive allocation}\left\{B_{i}=B_{i}^{*}\right\}$
Based on the operations of concatenation and Top-K in the first and second lines of Algorithm 2, we can flatten the second $\sum$ in $D^{\prime\prime}$ as follow:
$D^{\prime\prime}=2hC-2C\sum^{i\in[h],j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A,B)}A_{i}^{j}$ (35)
$\text{where}\>B=\sum^{i}B_{i}^{*}=\sum^{i}B_{i}$
Considering that $B=\sum{B_{i}}=\sum{B_{i}^{*}}$ , it is evident that $\sum^{i\in[h],j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A,B)}A_{i}^{j}\geq\sum^{i\in[h]}\sum^{j\in[n]}_{A_{i}^{j}\in\text{Top-K}(A_{i},B_{i})}A_{i}^{j}$. This is because under the premise of identical total budget, the global Top-K sum is greater than or equal to the sum of local Top-K sums of each head. Thus:
$D^{\prime\prime}\leq D^{\prime}$ (36)
Single-Doc. QA Multi-Doc. QA Summarization Few-shotLearning Synthetic Code
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Ave. Score
Full Cache 18.00 25.80 43.10 23.40 16.70 9.70 27.20 25.00 24.70 70.50 61.60 39.60 3.00 6.50 42.20 41.60  29.91
B=128h
H2O 17.90 17.73 36.10 21.52 17.51 9.26 16.13 22.99 19.64 43.50 60.64 36.36 3.00 5.50 34.93 36.74  24.97
StreamingLLM 12.81 11.32 29.04 17.24 13.67 6.91 16.34 20.25 17.35 41.00 52.74 25.77 0.50 3.00 28.38 30.98  20.46
SnapKV 17.51 17.57 38.89 22.15 17.28 9.13 15.01 21.96 17.94 46.00 61.05 35.97 0.00 4.00 36.92 37.83  24.95
Pyramid 18.17 17.58 39.08 22.05 16.78 8.13 14.74 22.24 17.88 47.50 60.11 37.02 0.50 3.50 36.96 38.73  25.06
Ada-SnapKV 18.64 18.61 39.59 22.51 17.05 9.19 15.28 22.88 18.98 52.50 61.69 36.76 0.00 3.00 36.82 39.63  25.82
Ada-Pyramid 18.35 18.93 39.49 22.57 16.83 8.61 15.05 23.22 18.85 55.50 60.93 37.39 0.50 3.50 36.55 39.79 26.00
B=256h
H2O 18.99 19.18 38.78 21.88 17.33 9.16 16.88 23.29 20.51 48.00 60.36 38.07 3.00 5.50 37.18 37.94  26.00
StreamingLLM 13.59 11.81 29.73 18.59 14.37 6.72 21.06 20.78 21.29 51.50 51.92 26.51 0.50 3.00 28.97 31.09  21.96
SnapKV 19.27 20.61 40.78 22.81 16.83 9.89 16.23 23.17 20.07 53.50 61.75 38.41 0.00 4.00 38.25 40.57  26.63
Pyramid 18.81 19.83 40.71 22.34 17.10 9.08 16.10 22.93 19.50 60.00 61.01 38.65 0.50 5.00 38.23 39.13  26.81
Ada-SnapKV 18.99 21.08 41.18 22.89 17.64 9.52 16.71 23.05 20.48 67.00 61.27 38.74 0.00 3.50 39.60 40.96  27.66
Ada-Pyramid 18.78 20.32 40.50 22.73 17.01 9.37 16.05 23.60 19.93 69.00 61.43 39.07 2.00 5.00 38.40 40.08 27.70
B=512h
H2O 18.61 20.07 39.82 22.08 17.21 10.13 17.62 23.65 21.41 54.50 61.84 38.74 3.00 5.50 39.23 40.08  27.09
StreamingLLM 13.94 13.13 33.06 18.26 14.44 7.41 25.24 21.00 23.78 60.50 52.04 26.31 1.00 3.00 30.10 31.76  23.44
SnapKV 18.45 21.96 42.01 23.25 17.42 9.88 17.68 23.62 21.30 68.00 61.77 39.02 1.00 4.50 40.09 40.79  28.17
Pyramid 18.46 22.85 42.24 23.27 16.75 9.45 17.41 24.62 21.20 70.00 60.61 39.32 3.00 6.50 39.63 40.78 28.51
Ada-SnapKV 18.83 22.39 42.15 23.52 18.27 9.63 17.66 23.99 21.23 70.00 61.72 38.93 2.00 4.50 40.11 41.28 28.51
Ada-Pyramid 18.64 22.86 41.81 23.61 16.67 9.45 17.35 23.75 20.79 70.00 60.66 39.61 3.00 5.50 39.87 40.99  28.41
B=1024h
H2O 17.11 22.34 41.26 22.09 17.47 9.60 18.82 23.94 22.49 61.00 62.33 38.68 3.00 5.50 41.23 41.18  28.00
StreamingLLM 14.78 16.77 37.64 18.77 14.63 7.39 26.43 21.47 24.21 67.00 53.00 25.99 0.50 3.00 31.51 32.31  24.71
SnapKV 18.45 24.18 42.50 23.53 17.32 10.23 19.00 24.26 23.04 69.50 62.22 39.88 3.00 5.50 41.15 41.91 29.10
Pyramid 18.48 24.87 42.11 23.45 16.97 9.84 18.93 24.50 22.77 69.50 61.65 39.73 2.50 5.00 41.07 41.27  28.91
Ada-SnapKV 18.94 23.68 43.27 23.28 17.15 9.89 18.58 23.46 22.65 70.00 62.24 39.83 2.50 5.50 41.68 42.88 29.10
Ada-Pyramid 19.00 23.83 43.36 23.48 17.03 9.32 18.70 24.11 22.61 69.50 61.83 39.75 2.50 6.00 40.85 41.80  28.98
Table 2: Comparison Based on LWM-Text-Chat-1M Among 16 Datasets
A.4 Computational Efficiency
We assess the computational efficiency of Ada-SnapKV and Ada-Pyramid under adaptive allocation, as well as the original versions of SnapKV and Pyramid, using a budget of 128h across various context lengths within needle-in-a-haystack test. As shown in Figure 8, the peak memory footprint during inference for Ada-SnapKV and Ada-Pyramid, as well as SnapKV and Pyramid, remains almost consistent as sequence length increases, significantly lower than that of vanilla Full Cache. Consequently, this allows the original sequence length of 37K to be extended to most 429K, achieving a 10.59-fold improvement. In terms of speed, as shown in 8, the decoding latency of the four strategies remains almost consistent and is independent of the context length, which is significantly lower than the decoding latency under Full Cache. This is primarily due to cache eviction, which greatly reduces the size of the KV Cache, thereby significantly alleviating the IO latency bottleneck in the autoregressive decoding stage.
Figure 7: Peak Memory Footprint
Figure 8: Decoding Latency
A.5 Detailed results for LWM model Among 16 Datasets
The table 2 presents quality scores of different eviction strategies based on the LWM model across 16 datasets. Overall, the results are consistent with those of Mistral, and the adaptive allocation also leads to quality improvements after cache eviction.
A.6 Detailed Information of Datasets
Table 3 provides a comprehensive description of information pertaining to 16 datasets.
A.7 Detailed Visualization of Head Concentration
Figure 9 supplements Figure 3 in the main paper by presenting the visualization results across all layers. It can be observed that in all layers, different heads exhibit significant variations in attention concentration. This indicates that the adaptive allocation algorithm has great potential to reduce the eviction loss in practice.
A.8 Code
We are in the process of making our source code available to the public.
Label Task Task Type Eval metric Avg len Language Sample Num
NrtvQA NarrativeQA Single-Doc. QA F1 18,409 EN 200
Qasper Qasper Single-Doc. QA F1 3,619 EN 200
MF-en MultiFieldQA-en Single-Doc. QA F1 4,559 EN 150
HotpotQA HotpotQA Multi-Doc. QA F1 9,151 EN 200
2WikiMQA 2WikiMultihopQA Multi-Doc. QA F1 4,887 EN 200
Musique MuSiQue Multi-Doc. QA F1 11,214 EN 200
GovReport GovReport Summarization Rouge-L 8,734 EN 200
QMSum QMSum Summarization Rouge-L 10,614 EN 200
MultiNews MultiNews Summarization Rouge-L 2,113 EN 200
TREC TREC Few-shotLearning Accuracy 5,177 EN 200
TriviaQA TriviaQA Few-shotLearning F1 8,209 EN 200
SAMSum SAMSum Few-shotLearning Rouge-L 6,258 EN 200
PCount PassageCount Synthetic Accuracy 11,141 EN 200
PRe PassageRetrieval-en Synthetic Accuracy 9,289 EN 200
Lcc LCC Code Edit Sim 1,235 Python/C#/Java 500
RB-P RepoBench-P Code Edit Sim 4,206 Python/Java 500
Table 3: Details of 16 Datasets
Figure 9: Visualization of Heads’ Concentrations