### v3

# Ada-KV: Optimizing KV Cache Eviction by Adaptive Budget Allocation for Efficient LLM Inference

Abstract
Large Language Models have excelled in various fields but encounter challenges in memory and time efficiency due to the expanding Key-Value (KV) cache required for long-sequence inference. Recent efforts try to reduce KV cache size to a given memory budget by evicting vast non-critical cache elements during runtime, while preserving generation quality. Our revisiting of current eviction methods reveals that they fundamentally minimize an upper bound of the eviction loss between the pre- and post-eviction outputs of multi-head self-attention mechanisms. Moreover, our analysis indicates that the common practices of uniformly assigning budgets across attention heads harm their post-eviction generation quality. In light of these findings, we propose a simple yet effective adaptive budget allocation algorithm. This algorithm not only optimizes the theoretical loss upper bound but also reduces the eviction loss in practice by aligning with the varied characteristics across different heads. By integrating this algorithm into two state-of-the-art methods, we demonstrate the effectiveness of using adaptive budget allocation to optimize KV cache eviction. Extensive evaluations on 16 datasets and the Needle-in-a-Haystack test confirm significant performance improvements across various tasks. Our code is available at https://github.com/FFY0/AdaKV.
1 Introduction
Autoregressive Large language models (LLMs) have achieved significant success and are widely utilized across diverse natural language processing applications, including dialogue systems (Yi et al. 2024), document summarization (Laban et al. 2023), and code generation (Gu 2023). The widespread deployments of LLMs have propelled the development of their capacities to process extended sequences. For instance, GPT-4 supports sequences up to 128K (Achiam et al. 2023), Claude3 up to 200K (Anthropic 2024), and Gemini-Pro-1.5 (Reid et al. 2024) up to 1M tokens. Modern LLMs are typically built with multiple layers of transformer blocks, each containing a multi-head self-attention layer. The computational cost of this self-attention mechanism increases quadratically with the token number of input sequence. Consequently, LLMs face significant efficiency challenges when processing long-sequence inputs.
For each multi-head self-attention layer, the inference process consists of two phases: prefilling and decoding. During prefilling, LLMs compute and store all Key-Value (KV) cache elements for the input tokens from the prompt. In subsequent decoding, LLMs autoregressively use the last token to retrieve information from past KV cache elements to generate each output token, continuing until the maximum length is reached or the process is actively stopped. However, when processing long-sequence inputs, the generated cache size can significantly increase, even far exceeding the model’s parameter size (Sun et al. 2024). This escalation leads to substantial memory challenges during both prefilling and decoding phases. Additionally, the extensive KV cache I/O during decoding incurs significant latency. This latency thus becomes a bottleneck of decoding, even surpassing the computation time (Tang et al. 2024).
To address the challenges posed by large KV cache sizes, various cache eviction methods have been developed, as highlighted in recent literature (Ge et al. 2023; Zhang et al. 2024b; Yang et al. 2024; Zhang et al. 2024a; Li et al. 2024). As depicted in Figure 1, these methods retain only the budgeted size of cache elements in each head, while evicting the others following the prefilling phase. Thus, they reduce the memory burden and enhance decoding speed, thereby facilitating efficient long-sequence inference. These eviction methods are designed with plug-and-play capabilities, allowing for straightforward integration into any LLM without the need for fine-tuning. They typically employ a Top-k selection strategy based on attention weights, to effectively distinguish between critical and non-critical cache elements, deciding which to retain and which to evict. Despite these advancements, the challenge of minimizing quality loss while employing these eviction methods remains unresolved in this field.
Our study begins by revisiting current Top-k eviction methods to uncover their underlying principles from a theoretical perspective. We reveal that they are equivalent to minimizing an upper bound of the eviction loss, which is quantified as the distance between the pre- and post-eviction outputs of self-attention mechanisms. Moreover, we find that the common practice of uniform budget allocation (Ge et al. 2023; Zhang et al. 2024b; Yang et al. 2024; Zhang et al. 2024a; Li et al. 2024) across different attention heads leads to a misallocation of resources, limiting the effectiveness of existing methods. Based on these findings, we propose a simple yet effective adaptive allocation algorithm. As an illustrative example shown in Figure 2, it effectively improves budget utilization by adaptively allocating overall budget across different attention heads based on their varied concentration degrees, thereby improving post-eviction generation quality. We further demonstrate the advantages of adaptive budget allocation both theoretically and empirically. Theoretically, it reduces the upper bound of eviction loss compared to uniform allocation. Empirically, its alignment with the varied concentration degrees among heads effectively reduces eviction loss in practice.
By integrating this adaptive allocation algorithm into two state-of-the-art (SOTA) methods, SnapKV (Li et al. 2024) and PyramidKV (Yang et al. 2024; Zhang et al. 2024a), we develop two adaptive eviction methods, Ada-SnapKV and Ada-Pyramid, respectively. Extensive evaluations across 16 standard datasets, covering various tasks in LongBench (Bai et al. 2023), demonstrate that both Ada-SnapKV and Ada-Pyramid significantly improve the generation quality. Additionally, the Needle-in-a-Haystack benchmark further confirms that adaptive allocation improves in-context retrieval ability. The main contributions are summarized as follows:
By defining eviction loss as the distance between pre- and post-eviction outputs, we reveal that current eviction methods equivalently minimize its theoretical upper bound within allocated budgets. We further identify that the prevalent uniform budget allocation in cache eviction impedes this minimization.
We introduce the first adaptive budget allocation algorithm for cache eviction, which both optimizes the theoretical upper bound and the practical eviction loss. By integrating this algorithm into two SOTA methods, we develop Ada-SnapKV and Ada-Pyramid, achieving significant enhancements as confirmed by comprehensive evaluations.
We provide a theoretical framework for analyzing cache eviction based on the upper bound of eviction loss and demonstrate practical advancements through efficient CUDA implementation. This framework and implementation pave the way for further advancements in optimizing cache eviction through adaptive allocation strategies.
2 Related Works
In the long-sequence inference, the vast scale of the KV cache elements leads to a memory-bound situation, causing significant memory burden and I/O latency (Wang and Chen 2023). Numerous studies have sought to mitigate this issue by reducing the cache size, notably through the eviction of non-critical cache elements111For additional related works, see Appendix A.1. These cache eviction methods are primarily divided into two categories: sliding window eviction and Top-k eviction methods. The sliding window eviction methods (2020; 2024; 2023), exemplified by StreamingLLM (Xiao et al. 2023), simply retain several initial cache elements and those within a sliding window, while evicting others. However, the undiscriminating sliding eviction of cache elements results in a significant reduction in generation quality. In contrast, Top-k eviction methods (2023; 2024b; 2024; 2024b; 2024; 2024a; 2024) identify and retain a selected set of critical cache elements based on attention weights, for enhancing the post-eviction generation quality. The adaptive budget allocation presented in this paper, tailored for Top-k Eviction Methods, enhances post-eviction generation quality within the same overall budget.
In the realm of Top-k eviction methods, early work like FastGen (Ge et al. 2023) searches and combines multiple strategies, such as maintaining caches of special elements, punctuation elements, recent elements, and Top-k selected elements, based on the characteristics of attention heads. H2O, as the representation of works (Zhang et al. 2024b; Liu et al. 2024b; Ren and Zhu 2024), develops a Top-k based eviction scheme that leverages the query states of all tokens to identify critical cache elements. However, due to the unidirectional attention mechanism, recent tokens accumulate fewer observations in these methods, leading to erroneous evictions. Recent works, such as SnapKV (Li et al. 2024) and Pyramid (Yang et al. 2024; Zhang et al. 2024a), address this issue by using query states within an observation window to identify critical elements, thereby achieving the SOTA performance. However, to our best knowledge, existing Top-k eviction methods typically assign the overall budget uniformly across different heads, resulting in misallocation. In contrast, our adaptive allocation algorithm, which has demonstrated superior theoretical and empirical results, provides a novel approach to optimizing existing methods.
3 Method
In this section, we begin by providing a formal description of a multi-head self-attention layer (Section 3.1). Building on this, we theoretically revisit the foundational principles of existing Top-k eviction methods by introducing an eviction loss metric (Section 3.2). Inspired by theoretical findings, we propose a simple yet effective algorithm for adaptive budget allocation, which is proven to outperform traditional uniform budget allocation, both theoretically and practically (Section 3.3). We further validate its compatibility with existing Top-k eviction methods by integrating it into two SOTA works, thereby improving their post-eviction generation quality (Section 3.4).
3.1 Preliminaries
LLMs are characterized by an autoregressive generation mode, where each step involves using the last token to predict the next token. Define as the embedding matrix of all tokens in the sequence, and as the last token used as input at the current time step. To clarify the subsequent theoretical exposition, we adopt the notation system from (Liu et al. 2023) under the assumption of attention heads, providing a formal description of one multi-head self-attention layer. The transformation matrices for each head , , , , map token embeddings to their respective Query, Key, and Value and the final output matrix transforms the intermediate result to the output hidden states. At each time step, the previous KV cache elements on head have been initialized as:
Then, the input token is mapped to its corresponding {query, key, value} for each head, and the previous KV cache is updated accordingly:
Finally, the output is computed using the attention weights as follows222The scaling factor is omitted for simplification.:
3.2 Revisiting the Top-k Cache Eviction
A set of indicator variables 333Given that the first dimension of is 1, is used to simplify the notation for . Similarly, is in the same manner. represent the eviction decision with allocated budgets for all heads :
where each element indicates whether the KV cache element in is evicted for head . Thus, only a budget size of cache elements is retained for head : and the overall budget for one attention layer is . Then, we can obtain the post-eviction output of multi-head self-attention mechanism:
where denotes element-wise multiplication. Theorem 1 further simplifies the post-eviction output by representing in terms of . The detailed proof is provided in Appendix A.3.
Theorem 1.
Given allocated budgets , the post-eviction output can be rewritten as:
The degradation of generation quality after cache eviction stems from changes in the attention output. Thus, we quantify the eviction loss as the distance between the pre- and post-eviction outputs of the self-attention mechanisms:
Utilizing the row norm of the matrix, we derive an upper bound for the Eviction Loss in Theorem 2. For a detailed proof, refer to Appendix A.4.
Theorem 2.
Given allocated budgets , the eviction loss caused by cache eviction can be bounded by :
where is a constant number, representing the max row norm among all matrices.
Top-k eviction methods typically presuppose the stability of critical cache elements during future generation process to facilitate cache eviction (2024b; 2024; 2024; 2024a). The SOTA methods (Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a) leverage query states of several tokens in an observation window to calculate observed attention weights with past KV cache elements, which, in conjunction with Top-k selections, approximate the identification of cache elements critical in subsequent generations. For simplicity, we assume that the window size is 1, implying the eviction procedure relies solely on a single query state associated with the last token for critical cache detection. As shown in Algorithm 1, Top-k selection only retains cache elements corresponding to the highest observed weights in each head , while evicting others. Obviously, given any allocated budgets , the eviction decision of Top-k selection maximizes the following equation:
Therefore, we establish Theorem 3, which demonstrates that the principle of Top-k eviction aims to minimize the upper bound of the eviction loss.
Input: Allocated Budgets , Observed Attention Weights associated with
Output: Top-k Eviction Decision
Theorem 3.
Given a budget allocation result , the Top-k cache eviction methods fundamentally minimize the upper bound of eviction loss:
3.3 Optimizing with Adaptive Budget Allocation
Theorems 2 and 3 demonstrate the minimization target depends on the eviction decision , thus is indirectly influenced by the budget allocation . This suggests that an appropriate allocation can further optimize compared to simple uniform allocation, which overlooks the distinct characteristics of each head in the self-attention mechanism—a discrepancy noted in other fields (Voita et al. 2019; Michel, Levy, and Neubig 2019; Clark et al. 2019). To address this issue, we propose the first adaptive budget allocation algorithm among heads for Top-k eviction. As outlined in Algorithm 2, it initially selects the largest observed attention weights from all heads within one layer. The frequency of selection for each head then informs the adaptive budget allocation , thereby optimizing the Top-k selection results in eviction procedure.
Theoretical Perspective: Adaptive vs. Uniform Allocation
The allocation results reshape the selection results of Algorithm 1, impacting the theoretical upper bound of eviction losses. Under the uniform budget allocation , the revised upper bound for is calculated as follows:
Conversely, under the adaptive budget allocation , its upper bound with is given by:
Theorem 4.
The adaptive budget allocation ensures its is consistently equal to or lower than with uniform budget allocation.
Input: Total Budget , Observed Attention Weights ;
Output: Allocated Budgets
Theorem 4 supports the theoretical advantage of adaptive budget allocation, which achieves an optimized eviction loss upper bound. The proof is provided in Appendix A.5.
Empirical Perspective: Adaptive vs. Uniform Allocation
Empirical evidence further substantiates that adaptive budget allocation capitalizes on the varied degrees of attention concentration among heads. Utilizing this variability is crucial for optimizing budget efficiency and minimizing eviction losses in practical scenarios. Figure 3 demonstrates that heads with sparse concentrations require significantly smaller budget proportions to achieve near-optimal top weight aggregation compared to more dispersed heads. Under such circumstances, the previous uniform budget allocation encounters a dilemma: it either wastes excessive and unwarranted budgets on heads with sparse concentrations or endures substantial eviction losses in dispersed heads. This significantly undermines the trade-off performance between the overall budget and post-eviction generation quality. In contrast, the adaptive budget allocation algorithm assigns large budgets to dispersed heads, while controlling the budget sizes for other sparse heads, effectively maintaining the overall budget size and mitigating the decline in generation quality. Figure 4 further illustrates that adaptive budget allocation consistently lowers practical eviction loss across all samples, underscoring its effectiveness in optimizing budget allocation.
3.4 Implementation of Seamless Integration
We demonstrate the seamless compatibility of our adaptive budget allocation algorithm with two SOTA methods, SnapKV and Pyramid, by integrating it into their existing Top-k eviction frameworks to create enhanced versions: Ada-SnapKV and Ada-Pyramid. Both SnapKV and Pyramid utilize tokens from a recent observation window (typically size 32) to identify and evict the less crucial elements in past KV cache. SnapKV excels under larger budget scenarios, while Pyramid is optimized for constrained budget conditions. This is because Pyramid employs a pyramidal form of budget distribution across different attention layers through pre-set hyper-parameters, favoring shallower layers, whereas SnapKV uniformly distributes the budget. Thus, for a specific layer with an overall budget of , their eviction algorithms are the same. However, both methods traditionally allocate budgets uniformly across all heads within one layer.
Incorporating our adaptive allocation, as outlined in Algorithm 3444For simplicity, Algorithm 3 presents the process sequentially; however, eviction operations can readily be parallelized in practice., we modify these methods to better manage budget allocation at the head level. This integration occurs prior to the eviction process in each layer, where our algorithm adaptively adjusts budget allocations based on the observed attention weights among heads, as shown in Line 6. Overall, they first calculate the observed attention weights of past cache elements using the query states within the observation window. A max pooling layer processes these weights to preserve essential information (Li et al. 2024), followed by a Top-k selection of past cache elements outside the observation window. These selected elements, along with others within the observation window, are retained, while the rest are evicted to reduce cache size. Moreover, we introduce a safeguard hyper-parameter, (defaulted to 0.5), to prevent the allocation of excessively small budgets to highly sparse heads, thereby enhancing fault tolerance for the presupposed stability of the critical elements (Li et al. 2024; Zhang et al. 2024b). To facilitate efficient handling of adaptively allocated, variable-number cache elements, we implement a flattened storage architecture using a custom CUDA kernel. This approach, combined with the Flash Attention technique (Kwon et al. 2023), ensures that the computational efficiency of Ada-SnapKV and Ada-Pyramid aligns with traditional methods. Further details on computational efficiency are provided in Appendix A.2.
| Single-Doc. QA | Multi-Doc. QA | Summarization | Few-shotLearning | Synthetic | Code | ||||||||||||
NrtvQA |
Qasper |
MF-en |
HotpotQA |
2WikiMQA |
Musique |
GovReport |
QMSum |
MultiNews |
TREC |
TriviaQA |
SAMSum |
PCount |
PRe |
Lcc |
RB-P |
Ave. Score |
| Full Cache | 26.63 | 32.99 | 49.34 | 42.77 | 27.35 | 18.77 | 32.87 | 24.24 | 27.10 | 71.00 | 86.23 | 42.96 | 2.75 | 86.98 | 55.33 | 52.87 | 42.51 |
| B=128h | |||||||||||||||||
| H2O | 21.19 | 21.66 | 38.60 | 30.63 | 20.65 | 12.19 | 20.65 | 22.42 | 21.81 | 39.00 | 82.52 | 40.68 | 2.98 | 79.56 | 49.13 | 46.76 | 34.40 |
| StreamingLLM | 16.61 | 14.74 | 31.40 | 28.05 | 21.36 | 12.08 | 18.44 | 18.91 | 19.26 | 43.50 | 74.22 | 29.00 | 2.75 | 31.65 | 41.27 | 38.84 | 27.63 |
| SnapKV | 19.17 | 21.40 | 42.93 | 36.76 | 22.44 | 15.86 | 19.16 | 21.84 | 21.55 | 47.50 | 84.15 | 40.24 | 2.30 | 68.26 | 50.69 | 47.13 | 35.09 |
| Pyramid | 20.16 | 21.77 | 43.55 | 36.78 | 23.12 | 14.39 | 19.53 | 22.03 | 21.47 | 51.00 | 84.62 | 40.24 | 2.79 | 70.77 | 50.57 | 46.53 | 35.58 |
| Ada-SnapKV | 20.63 | 22.58 | 45.68 | 37.90 | 23.49 | 16.55 | 19.99 | 22.28 | 21.55 | 59.50 | 85.00 | 40.62 | 3.09 | 69.36 | 50.98 | 48.17 | 36.71 |
| Ada-Pyramid | 20.50 | 21.71 | 45.61 | 36.81 | 23.57 | 15.84 | 19.75 | 22.13 | 22.00 | 60.50 | 84.04 | 40.51 | 3.21 | 73.60 | 51.24 | 48.02 | 36.81 |
| B=256h | |||||||||||||||||
| H2O | 21.54 | 22.92 | 42.56 | 31.07 | 22.53 | 13.76 | 22.52 | 22.40 | 23.09 | 40.50 | 84.20 | 40.77 | 3.41 | 86.10 | 50.98 | 48.17 | 36.03 |
| StreamingLLM | 17.93 | 16.01 | 33.36 | 30.71 | 21.30 | 10.08 | 20.66 | 19.47 | 22.89 | 53.50 | 73.59 | 29.22 | 3.00 | 27.77 | 42.30 | 39.87 | 28.85 |
| SnapKV | 22.37 | 23.74 | 48.13 | 38.56 | 22.43 | 15.66 | 21.91 | 23.13 | 23.15 | 61.50 | 85.45 | 41.42 | 3.09 | 84.54 | 53.22 | 50.24 | 38.66 |
| Pyramid | 20.09 | 24.00 | 47.33 | 38.24 | 22.48 | 16.02 | 21.40 | 22.45 | 22.63 | 63.00 | 84.93 | 40.98 | 3.40 | 82.48 | 52.78 | 49.36 | 38.22 |
| Ada-SnapKV | 22.55 | 25.78 | 48.33 | 40.30 | 24.24 | 16.64 | 21.63 | 23.03 | 23.19 | 67.00 | 85.78 | 41.53 | 3.47 | 87.07 | 53.86 | 51.13 | 39.72 |
| Ada-Pyramid | 22.64 | 24.64 | 47.40 | 40.25 | 23.62 | 16.83 | 21.82 | 23.34 | 22.70 | 66.50 | 84.99 | 41.34 | 2.78 | 86.90 | 53.17 | 49.52 | 39.28 |
| B=512h | |||||||||||||||||
| H2O | 21.72 | 26.03 | 44.81 | 32.33 | 23.16 | 14.86 | 23.65 | 22.84 | 24.70 | 42.00 | 85.22 | 41.57 | 3.40 | 86.45 | 53.04 | 49.68 | 37.22 |
| StreamingLLM | 18.76 | 17.17 | 37.09 | 30.21 | 21.64 | 9.93 | 24.44 | 20.00 | 25.57 | 62.00 | 72.36 | 29.95 | 2.48 | 18.17 | 43.70 | 40.13 | 29.60 |
| SnapKV | 24.60 | 27.81 | 48.98 | 39.46 | 25.25 | 16.98 | 23.70 | 22.96 | 24.37 | 67.00 | 85.88 | 41.26 | 2.78 | 86.56 | 54.81 | 51.71 | 40.26 |
| Pyramid | 23.23 | 27.94 | 48.87 | 40.50 | 24.36 | 16.74 | 23.22 | 23.16 | 24.37 | 67.00 | 85.73 | 41.74 | 3.16 | 85.67 | 54.16 | 50.34 | 40.01 |
| Ada-SnapKV | 23.39 | 28.72 | 48.96 | 40.60 | 25.20 | 17.25 | 23.15 | 23.48 | 24.41 | 68.00 | 86.39 | 41.69 | 2.73 | 88.92 | 54.69 | 51.51 | 40.57 |
| Ada-Pyramid | 24.03 | 28.98 | 48.39 | 39.25 | 24.50 | 18.38 | 23.13 | 23.90 | 24.30 | 68.00 | 85.89 | 41.89 | 2.98 | 87.71 | 54.46 | 51.39 | 40.45 |
| B=1024h | |||||||||||||||||
| H2O | 23.90 | 28.62 | 46.46 | 37.03 | 24.74 | 15.04 | 25.30 | 23.11 | 25.92 | 46.00 | 85.93 | 41.80 | 3.24 | 86.57 | 54.46 | 51.01 | 38.70 |
| StreamingLLM | 19.42 | 21.69 | 41.75 | 32.40 | 22.18 | 11.18 | 27.13 | 21.09 | 26.59 | 67.00 | 71.79 | 30.11 | 2.88 | 16.57 | 44.82 | 39.76 | 31.02 |
| SnapKV | 25.47 | 29.57 | 49.33 | 40.90 | 25.53 | 19.01 | 25.94 | 23.89 | 26.21 | 69.50 | 86.48 | 42.10 | 2.98 | 88.56 | 55.57 | 51.92 | 41.44 |
| Pyramid | 24.21 | 29.86 | 48.93 | 40.75 | 25.05 | 18.77 | 25.73 | 24.06 | 25.65 | 68.50 | 86.31 | 42.25 | 2.97 | 87.17 | 54.75 | 52.10 | 41.07 |
| Ada-SnapKV | 24.79 | 31.94 | 48.45 | 40.73 | 26.22 | 19.11 | 25.61 | 23.92 | 26.03 | 70.00 | 86.32 | 42.35 | 2.91 | 88.31 | 55.44 | 52.55 | 41.54 |
| Ada-Pyramid | 25.09 | 30.94 | 48.18 | 40.00 | 26.52 | 19.10 | 24.93 | 23.71 | 25.86 | 70.00 | 86.34 | 42.64 | 2.56 | 86.92 | 54.93 | 51.90 | 41.23 |
4 Experiments
4.1 Settings
Datasets
Firstly, we carry out a comprehensive evaluation using 16 datasets within LongBench (Bai et al. 2023), a long-sequence benchmark covering multi-task domains of single-document QA (Kočiskỳ et al. 2018; Dasigi et al. 2021), multi-document QA (Yang et al. 2018; Ho et al. 2020; Trivedi et al. 2022), summarization (Huang et al. 2021; Zhong et al. 2021; Fabbri et al. 2019), few-shot learning (Joshi et al. 2017; Gliwa et al. 2019; Li and Roth 2002), synthetic tasks (Bai et al. 2023), and code generation (Guo et al. 2023; Liu, Xu, and McAuley 2023). These datasets feature varying average input lengths from 1,235 to 18,409 tokens, necessitating extensive KV cache size for inference, and thereby rendering them suitable for evaluating KV cache eviction methods under different memory budgets. Each dataset is assessed using LongBench-recommended metrics, with each quality scores up to 100. Detailed dataset information is provided in Appendix A.6. Additionally, we also employ the widely-used ’Needle-in-a-Haystack’ test to specifically examine the impact of proposed adaptive budget allocation on the models’ fundamental long-context retrieval capabilities.
Baselines
We select the SnapKV(Li et al. 2024) and Pyramid(Yang et al. 2024; Zhang et al. 2024a) as the primary baselines, given that they are leading approaches and foundational bases for our Ada-SnapKV and Ada-Pyramid methods Additionally, StreamingLLM (Xiao et al. 2023) represents Sliding Window Eviction Methods, and H2O (Zhang et al. 2024b) exemplifies early Top-K Eviction Methods based on all query states in our baseline comparisons.
Base Models
In the experiments, we employ two open-source base models: Mistral-7B-instruct-v0.2 (Jiang et al. 2023) and LWM-Text-Chat-1M (Liu et al. 2024a). The Mistral 7B model features a context length of 32K and has been adopted as the primary model in related studies (Li et al. 2024; Zhang et al. 2024a) due to its moderate parameter size and remarkable capability for long-sequence tasks. Meanwhile, the LWM 7B model stands as the state of the art with its 1M context length, facilitating evaluations under extreme context lengths in the Needle-in-a-Haystack test.
Parameters
Considering the minimum average length of all datasets is 1235, we assess all methods under four varied layer budget sizes of for comprehensive evaluations. We follow the common practice from prior studies (Li et al. 2024; Yang et al. 2024; Zhang et al. 2024a), conducting cache eviction methods after each layer’s prefilling phase for comparison. In all experiments, the hyper-parameter in adaptive budget allocation is set to 0.5. Both Ada-SnapKV and Ada-Pyramid, as well as SnapKV and Pyramid, utilize the same configuration settings as described in (Li et al. 2024), ensuring comparability with the observation window size of 32 and a max pooling kernel size of 7. Parameters for StreamingLLM and H2O conform to the default settings in (Zhang et al. 2024b; Xiao et al. 2023). All experiments are conducted on single A100-80G GPU. For more details, please refer to our code in supplementary materials.
4.2 Evaluations Among 16 Datasets
Detailed results for each dataset based on Mistral model are provided in Table 1 and other results based on LWM model are placed in Appendix A.7 due to space constraints. We take a budget as an example shown in Table 1 to demonstrate the improvements. By integrating the adaptive budget allocation, Ada-SnapKV enhances the quality scores in 15 out of 16 datasets compared to the original SnapKV, increasing the average score from 35.09 to 36.71. Similarly, Ada-Pyramid surpasses the original Pyramid in 14 of 16 datasets, boosting the average score from 35.58 to 36.81.
Figure 5 summarizes average scores of all methods based on Mistral and LWM across 16 datasets. Notably, StreamingLLM, as a representative of sliding window eviction methods, generates significantly lower quality outputs due to its inefficiency in identifying important cache elements. SnapKV and Pyramid, employing Top-k selection with an observation window, exhibit closely matched performance, surpassing previous H2O. Additionally, our Ada-SnapKV and Ada-Pyramid methods enhance generated quality across various budgets. The two Ada-enhanced methods alternately lead and surpass base versions, especially in small budgets. Such consistent improvement underscores the necessity and effectiveness of adaptive budget allocation, as supported by both theoretical derivations and empirical evidence.
4.3 Evaluations on Needle-in-a-Haystack Test
As shown in Figure 6, we employ a Needle-in-a-Haystack test to demonstrate the enhancement of long-context retrieval capabilities through adaptive budget allocation. Consistent with previous experiments, all configurations maintain a observation window size of 32 and a pooling kernel size of 7, with the maximum inference length limited to 37K for the full cache case on an A100-80G. Under a cache budget of , Ada-SnapKV and SnapKV extend the maximum length up to 429K, while the Ada-Pyramid and Pyramid extend to 365K. Significantly, both Ada-SnapKV and Ada-Pyramid improve long-text retrieval capabilities compared to previous SnapKV and Pyramid. In particular, Ada-SnapKV and Ada-Pyramid achieve near-lossless retrieval within the original 37K length, a feat not replicated by the standard SnapKV and Pyramid. In terms of average scores, Ada-SnapKV improves from 94.84 to 95.99, while Ada-Pyramid increases from 96.02 to 96.69. Additional evaluations of memory and time computational efficiency across varied lengths in the Needle-in-a-Haystack test, are available in Appendix A.2, demonstrating that the adaptive allocation preserves computational efficiency consistent with original version.
5 Conclusion
In this study, we revisit prevailing cache eviction methods for efficient long-sequence inference, revealing that they primarily minimize the upper bound of the distance between pre- and post-eviction outputs. Based on this insight, we propose the first adaptive budget allocation algorithm for optimizing the KV cache eviction, which theoretically lowers the upper bound compared to previous methods. Our empirical studies also indicate that this adaptive algorithm leverages the varying degrees of attention concentration within the multi-head self-attention mechanism. The development of two novel adaptive eviction methods, Ada-SnapKV and Ada-Pyramid, which incorporate this adaptive allocation, demonstrates remarkable improvements in comprehensive evaluations. Our work highlights the substantial potential for advancing cache eviction methods through our theoretical framework and adaptive budget allocation implementation, specifically designed to exploit the unique characteristics of different attention heads in LLMs.
Appendix A Appendix
A.1 Additional Related Works
Additional works also mitigate the challenges posed by massive KV Caches during long-sequence inference while not reducing the number of cache elements. These works are fundamentally orthogonal to our work. For instance, in our implementation, we have integrated the Flash Attention (Dao et al. 2022) technique to enhance efficient computation. Similar efforts, such as Page Attention (Kwon et al. 2023), employ efficient memory management strategies to reduce I/O latency without altering the size of the KV Cache. Other works, called KV cache quantization, reduce the size of cache by lowering the precision of individual elements. Our cache eviction techniques also be further combined and complemented with quantization in the future. In our experiments, we have employed 16-bit half-precision for inference. Under these conditions, further quantization only yields limited benefits. However, in our experiments, cache eviction methods are able to compress the cache size to below 10% with minor quality loss. A recent work (Tang et al. 2024) attempts to reduce I/O latency by only recalling KV cache elements relevant to the current query for computation. However, it is constrained by substantial memory burden, making deployment on GPUs with limited storage capacities challenging. Future efforts could further reduce memory overhead and decrease I/O latency by collaboratively employing cache eviction and recalling techniques.
A.2 Computational Efficiency
Cache eviction methods aim to improve the memory and time efficiency of LLM inference by evicting the vast KV cache elements to reduce the memory burden and enhance decoding speed. Thus we assess the peak memory footprint and decoding latency of Ada-SnapKV and Ada-Pyramid, along with the original SnapKV and Pyramid versions, across various context lengths in the Needle-in-a-Haystack test to demonstrate their consistent computational efficiency under the same budget. As shown in Figure 8, the peak memory footprint during inference for Ada-SnapKV and Ada-Pyramid, as well as SnapKV and Pyramid, remains the same as sequence length increases, significantly lower than that of vanilla Full Cache. Consequently, this allows the original sequence length of 37K to be extended to most 429K, achieving a 10.59-fold improvement. In terms of speed, as shown in 8, the decoding latency of the four strategies remains almost consistent and is independent of the context length, which is significantly lower than the decoding latency under Full Cache. This is primarily due to cache eviction, which greatly reduces the size of the KV Cache, thereby significantly alleviating the IO latency bottleneck in the autoregressive decoding phase.
| Label | Task | Task Type | Eval metric | Avg len | Language | Sample Num |
| NrtvQA | NarrativeQA | Single-Doc. QA | F1 | 18,409 | EN | 200 |
| Qasper | Qasper | Single-Doc. QA | F1 | 3,619 | EN | 200 |
| MF-en | MultiFieldQA-en | Single-Doc. QA | F1 | 4,559 | EN | 150 |
| HotpotQA | HotpotQA | Multi-Doc. QA | F1 | 9,151 | EN | 200 |
| 2WikiMQA | 2WikiMultihopQA | Multi-Doc. QA | F1 | 4,887 | EN | 200 |
| Musique | MuSiQue | Multi-Doc. QA | F1 | 11,214 | EN | 200 |
| GovReport | GovReport | Summarization | Rouge-L | 8,734 | EN | 200 |
| QMSum | QMSum | Summarization | Rouge-L | 10,614 | EN | 200 |
| MultiNews | MultiNews | Summarization | Rouge-L | 2,113 | EN | 200 |
| TREC | TREC | Few-shotLearning | Accuracy | 5,177 | EN | 200 |
| TriviaQA | TriviaQA | Few-shotLearning | F1 | 8,209 | EN | 200 |
| SAMSum | SAMSum | Few-shotLearning | Rouge-L | 6,258 | EN | 200 |
| PCount | PassageCount | Synthetic | Accuracy | 11,141 | EN | 200 |
| PRe | PassageRetrieval-en | Synthetic | Accuracy | 9,289 | EN | 200 |
| Lcc | LCC | Code | Edit Sim | 1,235 | Python/C#/Java | 500 |
| RB-P | RepoBench-P | Code | Edit Sim | 4,206 | Python/Java | 500 |
| Single-Doc. QA | Multi-Doc. QA | Summarization | Few-shotLearning | Synthetic | Code | ||||||||||||
NrtvQA |
Qasper |
MF-en |
HotpotQA |
2WikiMQA |
Musique |
GovReport |
QMSum |
MultiNews |
TREC |
TriviaQA |
SAMSum |
PCount |
PRe |
Lcc |
RB-P |
Ave. Score |
| Full Cache | 18.00 | 25.80 | 43.10 | 23.40 | 16.70 | 9.70 | 27.20 | 25.00 | 24.70 | 70.50 | 61.60 | 39.60 | 3.00 | 6.50 | 42.20 | 41.60 | 29.91 |
| B=128h | |||||||||||||||||
| H2O | 17.90 | 17.73 | 36.10 | 21.52 | 17.51 | 9.26 | 16.13 | 22.99 | 19.64 | 43.50 | 60.64 | 36.36 | 3.00 | 5.50 | 34.93 | 36.74 | 24.97 |
| StreamingLLM | 12.81 | 11.32 | 29.04 | 17.24 | 13.67 | 6.91 | 16.34 | 20.25 | 17.35 | 41.00 | 52.74 | 25.77 | 0.50 | 3.00 | 28.38 | 30.98 | 20.46 |
| SnapKV | 17.51 | 17.57 | 38.89 | 22.15 | 17.28 | 9.13 | 15.01 | 21.96 | 17.94 | 46.00 | 61.05 | 35.97 | 0.00 | 4.00 | 36.92 | 37.83 | 24.95 |
| Pyramid | 18.17 | 17.58 | 39.08 | 22.05 | 16.78 | 8.13 | 14.74 | 22.24 | 17.88 | 47.50 | 60.11 | 37.02 | 0.50 | 3.50 | 36.96 | 38.73 | 25.06 |
| Ada-SnapKV | 18.64 | 18.61 | 39.59 | 22.51 | 17.05 | 9.19 | 15.28 | 22.88 | 18.98 | 52.50 | 61.69 | 36.76 | 0.00 | 3.00 | 36.82 | 39.63 | 25.82 |
| Ada-Pyramid | 18.35 | 18.93 | 39.49 | 22.57 | 16.83 | 8.61 | 15.05 | 23.22 | 18.85 | 55.50 | 60.93 | 37.39 | 0.50 | 3.50 | 36.55 | 39.79 | 26.00 |
| B=256h | |||||||||||||||||
| H2O | 18.99 | 19.18 | 38.78 | 21.88 | 17.33 | 9.16 | 16.88 | 23.29 | 20.51 | 48.00 | 60.36 | 38.07 | 3.00 | 5.50 | 37.18 | 37.94 | 26.00 |
| StreamingLLM | 13.59 | 11.81 | 29.73 | 18.59 | 14.37 | 6.72 | 21.06 | 20.78 | 21.29 | 51.50 | 51.92 | 26.51 | 0.50 | 3.00 | 28.97 | 31.09 | 21.96 |
| SnapKV | 19.27 | 20.61 | 40.78 | 22.81 | 16.83 | 9.89 | 16.23 | 23.17 | 20.07 | 53.50 | 61.75 | 38.41 | 0.00 | 4.00 | 38.25 | 40.57 | 26.63 |
| Pyramid | 18.81 | 19.83 | 40.71 | 22.34 | 17.10 | 9.08 | 16.10 | 22.93 | 19.50 | 60.00 | 61.01 | 38.65 | 0.50 | 5.00 | 38.23 | 39.13 | 26.81 |
| Ada-SnapKV | 18.99 | 21.08 | 41.18 | 22.89 | 17.64 | 9.52 | 16.71 | 23.05 | 20.48 | 67.00 | 61.27 | 38.74 | 0.00 | 3.50 | 39.60 | 40.96 | 27.66 |
| Ada-Pyramid | 18.78 | 20.32 | 40.50 | 22.73 | 17.01 | 9.37 | 16.05 | 23.60 | 19.93 | 69.00 | 61.43 | 39.07 | 2.00 | 5.00 | 38.40 | 40.08 | 27.70 |
| B=512h | |||||||||||||||||
| H2O | 18.61 | 20.07 | 39.82 | 22.08 | 17.21 | 10.13 | 17.62 | 23.65 | 21.41 | 54.50 | 61.84 | 38.74 | 3.00 | 5.50 | 39.23 | 40.08 | 27.09 |
| StreamingLLM | 13.94 | 13.13 | 33.06 | 18.26 | 14.44 | 7.41 | 25.24 | 21.00 | 23.78 | 60.50 | 52.04 | 26.31 | 1.00 | 3.00 | 30.10 | 31.76 | 23.44 |
| SnapKV | 18.45 | 21.96 | 42.01 | 23.25 | 17.42 | 9.88 | 17.68 | 23.62 | 21.30 | 68.00 | 61.77 | 39.02 | 1.00 | 4.50 | 40.09 | 40.79 | 28.17 |
| Pyramid | 18.46 | 22.85 | 42.24 | 23.27 | 16.75 | 9.45 | 17.41 | 24.62 | 21.20 | 70.00 | 60.61 | 39.32 | 3.00 | 6.50 | 39.63 | 40.78 | 28.51 |
| Ada-SnapKV | 18.83 | 22.39 | 42.15 | 23.52 | 18.27 | 9.63 | 17.66 | 23.99 | 21.23 | 70.00 | 61.72 | 38.93 | 2.00 | 4.50 | 40.11 | 41.28 | 28.51 |
| Ada-Pyramid | 18.64 | 22.86 | 41.81 | 23.61 | 16.67 | 9.45 | 17.35 | 23.75 | 20.79 | 70.00 | 60.66 | 39.61 | 3.00 | 5.50 | 39.87 | 40.99 | 28.41 |
| B=1024h | |||||||||||||||||
| H2O | 17.11 | 22.34 | 41.26 | 22.09 | 17.47 | 9.60 | 18.82 | 23.94 | 22.49 | 61.00 | 62.33 | 38.68 | 3.00 | 5.50 | 41.23 | 41.18 | 28.00 |
| StreamingLLM | 14.78 | 16.77 | 37.64 | 18.77 | 14.63 | 7.39 | 26.43 | 21.47 | 24.21 | 67.00 | 53.00 | 25.99 | 0.50 | 3.00 | 31.51 | 32.31 | 24.71 |
| SnapKV | 18.45 | 24.18 | 42.50 | 23.53 | 17.32 | 10.23 | 19.00 | 24.26 | 23.04 | 69.50 | 62.22 | 39.88 | 3.00 | 5.50 | 41.15 | 41.91 | 29.10 |
| Pyramid | 18.48 | 24.87 | 42.11 | 23.45 | 16.97 | 9.84 | 18.93 | 24.50 | 22.77 | 69.50 | 61.65 | 39.73 | 2.50 | 5.00 | 41.07 | 41.27 | 28.91 |
| Ada-SnapKV | 18.94 | 23.68 | 43.27 | 23.28 | 17.15 | 9.89 | 18.58 | 23.46 | 22.65 | 70.00 | 62.24 | 39.83 | 2.50 | 5.50 | 41.68 | 42.88 | 29.10 |
| Ada-Pyramid | 19.00 | 23.83 | 43.36 | 23.48 | 17.03 | 9.32 | 18.70 | 24.11 | 22.61 | 69.50 | 61.83 | 39.75 | 2.50 | 6.00 | 40.85 | 41.80 | 28.98 |
A.3 Proof of Theorem 1
Theorem.
Given allocated budgets , the post-eviction output can be rewritten as:
Proof.
Consider the softmax function as:
Thus, the attention weight after eviction procedure is:
Given , we can get .
Then we can obtain:
Thus:
A.4 Proof of Theorem 2
Theorem.
Given allocated budgets , the eviction loss caused by cache eviction can be bounded by :
where is a constant number, representing the max row norm among all matrices.
Proof.
By calculating the L1 distance between their outputs, we can obtain
By expanding , we can further simplify the expression.
Considering
Finally,
A.5 Proof of Theorem 4
Theorem.
The adaptive budget allocation ensures its is consistently equal to or lower than with uniform budget allocation.
Proof.
Considering is the selection result of Top-k algorithm based on uniform budget allocation , it is evident that
This is because under the premise of identical total budget, the sum of global Top-k is greater than or equal to the sum of local Top-k sums of each head. Thus:
A.6 Detailed Information of Datasets
Table 2 provides a comprehensive description of information pertaining to 16 datasets in LongBench.
A.7 Detailed results for LWM model Among 16 Datasets
Table 3 presents quality scores of different eviction strategies based on the LWM model across 16 datasets. Overall, the results are consistent with those of Mistral, and the adaptive allocation also leads to quality improvements after cache eviction.
A.8 Detailed Visualization of Head Concentration
Figure 9 supplements Figure 3 in the main paper by presenting the visualization results across all layers. It can be observed that in all layers, different heads exhibit significant variations in attention concentration. This indicates that the adaptive allocation algorithm has great potential to reduce the eviction loss in practice.